### v1

# SonarSplat: Novel View Synthesis of Imaging Sonar via Gaussian Splatting

Abstract
In this paper, we present SonarSplat, a novel Gaussian splatting framework for imaging sonar that demonstrates realistic novel view synthesis and models acoustic streaking phenomena. Our method represents the scene as a set of 3D Gaussians with acoustic reflectance and saturation properties. We develop a novel method to efficiently rasterize learned Gaussians to produce a range/azimuth image that is faithful to the acoustic image formation model of imaging sonar. In particular, we develop a novel approach to model azimuth streaking in a Gaussian splatting framework. We evaluate SonarSplat using real-world datasets of sonar images collected from an underwater robotic platform in a controlled test tank and in a real-world river environment. Compared to the state-of-the-art, SonarSplat offers improved image synthesis capabilities (+2.5 dB PSNR). We also demonstrate that SonarSplat can be leveraged for azimuth streak removal and 3D scene reconstruction.
1 Introduction
Figure 1: We present SonarSplat, a method that uses posed sonar images to produce a novel 3D Gaussian Splat representation that enables novel view synthesis of realistic sonar images and coarse 3D reconstruction. With SonarSplat, the scene is represented with a set of 3D Gaussians that model the reflectivity and azimuth streaking probabilities of objects in the rendering process. This novel formulation enables additional features of SonarSplat, such as azimuth streak removal, enabling the recovery of clean sonar images from saturated ones.
Acoustic sensors, such as imaging sonar, are commonly used for infrastructure inspection, large-area mapping, and target detection in underwater environments [12, 34, 20]. Unlike optical sensors, which are severely range-limited due to water column effects on light propagation, acoustic sensors can capture data at long ranges to provide critical information about subsea environments. Although sonar exhibits many desirable qualities such as longer range, invariance to lighting conditions, and the ability to discern material properties, there exist acoustic phenomena, including elevation ambiguity, azimuth streaking, and multi-path reflections, which make sonar interpretation difficult for operators and computer vision algorithms alike. Furthermore, the severe lack of large, public datasets for sonar slows progress in traditional computer vision tasks like object detection, segmentation, and reconstruction.
Recently, neural radiance fields (NeRFs) [22] have demonstrated the potential for high-fidelity data synthesis and denoising of optical imagery [27, 6]. The benefits of neural rendering have been recognized by the underwater perception community, with prior work exploring neural rendering for 3D object reconstruction using underwater cameras [15, 35] and sonar data [30, 31, 44, 7, 43]. While these innovations are promising, training and processing NeRFs is extremely costly and time-consuming, making deploying them in real-time or on resource-constrained devices difficult. More recently, 3D Gaussian splatting (3DGS) [13] was developed as a faster alternative to NeRF. 3DGS utilizes sparse points in the form of 3D Gaussians to represent the scene. This allows 3DGS to preserve properties of continuous volumetric radiance fields while lowering computational costs. Gaussian splatting has been leveraged for underwater imagery [45, 16, 23]. Most relevant to our proposed work, ZSplat proposes a Gaussian splatting framework for RGB-sonar fusion [32]. However, ZSplat leverages the fusion of sonar data to improve the rendering of RGB images and does not focus on enabling high-fidelity data synthesis for sonar imagery or provide evaluation for the quality of rendered sonar images. Thus, there is a clear gap in the literature for a framework capable of efficient and effective sonar image synthesis.
In this paper, we propose SonarSplat (Fig. 1), a novel Gaussian splatting framework for imaging sonar that enables efficient and high quality novel view synthesis for imaging sonar data. We focus on modeling characteristic acoustic noise properties, specifically, azimuth streaking, or receiver saturation, which can make image interpretation and 3D reconstruction difficult. By modeling realistic noise effects, we demonstrate that we can leverage SonarSplat to perform efficient, high quality sonar image synthesis and noise removal. Lastly, we also demonstrate SonarSplat’s ability to learn geometry of the scene, overcoming the elevation ambiguity inherent in imaging sonar image formation. Our main contributions are summarized as follows:
  • We develop a novel 3D Gaussian splatting framework for rasterization of range/azimuth sonar images from known and novel viewpoints.
  • We develop a novel method for learning the probability of azimuth streaking in a differentiable manner, which allows for removal of unwanted azimuth streaking artifacts from rendered images.
  • We perform quantitative and qualitative evaluation of our method on real imaging sonar data collected from a robotic platform with a variety of sonar ranges and environments.
Code and data will be made publicly available on a project website after a double-blind review.
2 Related Work
Figure 2: Overview of SonarSplat. Our method takes as input a sensor pose and an initial set of 3D Gaussians representing the scene. Then, we transform the 3D Gaussians into the sensor’s image space. We splat the reflectivity parameter $\nu_{k}$ to get the unsaturated image $I_{u}$. Additionally, we optimize and splat per-Gaussian azimuth streaking probabilities $p_{k}$. All the probabilities in range interval $r_{i}$ are considered in our novel Adaptive Gain Module, which adjusts the receiver gain applied to $I_{u}$. Finally, we produce $\hat{I}$ by multiplying the gain $A$ by $I_{u}$. All parameters are optimized using gradient descent by taking losses with respect to the sonar image pixel values. All sonar images shown are polar (range-azimuth) coordinates.
2.1 Radiance Field Methods
NeRFs represent scenes implicitly using neural networks and have been explored for a variety of applications, including novel view synthesis, inverse rendering, data generation, and image denoising [21, 8, 27, 6]. Later works improved NeRFs through improved sampling of the volumetric representation and improvements in rendering speed with alternative encoding methods [24, 2]. Despite their vast success as a learned scene representation, a key struggle for the practical application of NeRFs is the long training time and the slow rendering speeds, primarily due to the computationally expensive volumetric rendering process [21]. A recent development in radiance field methods is Gaussian splatting, which uses rasterization instead of volumetric rendering to optimize and produce a scene representation [13]. The resulting renders from 3DGS are both high in quality and reach unprecedented rendering speeds, enabling fast rendering for various real-time applications [19, 18].
Despite their impressive performance and rapid adoption, radiance field methods were developed with optical sensors in mind, limiting their direct application to non-visual sensors. Recent developments have integrated the physics-based image formation models of various non-visual sensors, including radar, LIDAR, sonar, thermal cameras, and event cameras into NeRF and Gaussian splatting frameworks for various vision tasks [3, 4, 10, 30, 17]. In this work, we propose a novel Gaussian splatting framework for imaging sonar for underwater applications.
2.2 Perception Challenges for Imaging Sonar
Imaging sonar is a popular underwater sensor for perception tasks like object detection, segmentation, reconstruction, and inspection [30, 9, 12, 38, 25]. The key challenges with working with imaging sonars for these tasks are properties inherent with the acoustic image formation model [29]. The sensor is susceptible to multi-path effects, highly nonlinear noise patterns, and phenomena like azimuth streaking [28]. Azimuth streaking, in particular, can make sonar image understanding difficult by suppressing returns from certain azimuth angles, causing black bars in the sonar image [28]. In addition, the sensor inherently loses information from 3D into 2D when capturing the acoustic returns, causing an effect known as elevation ambiguity [40]. This ambiguity has made the successful application of imaging sonar to 3D vision tasks difficult [25, 11, 42]. Yet, imaging sonars still remain an extremely valuable sensor for underwater imaging, especially in long range imaging and deep water applications where optical sensors may fail [39].
Given its importance to underwater perception, prior work has proposed sonar simulators to validate perception algorithms without the cost, time commitment, and difficulty of real-world data collection [29, 5]. Although these works model complex interactions with the environment, such as multi-path reflections and azimuth streaking, there still exists a sim-to-real gap inherent with simulators. Furthermore, simulators also require the manual tuning of acoustic and scene parameters to match the distribution of data desired by the end-user. Rather, our proposed work proposes a Gaussian splatting framework that can learn from real data to produce realistic sonar data from novel poses, which may be useful for augmenting datasets and addressing the data scarcity problem.
2.3 Novel View Synthesis for Imaging Sonar
Prior work has incorporated the sonar image formation model into signed distance field (SDF) and NeRF formulations [30, 7]. Neusis [30] presents a NeRF framework for dense 3D reconstruction of objects using an imaging sonar, specifically, a forward-looking sonar (FLS). Neusis proposes a differentiable volumetric renderer to model the propagation of acoustic waves to synthesize imaging sonar measurements. By using this implicit representation and multiple sensor measurements, Neusis attempts to resolve the elevation ambiguity present in the image data to construct a 3D mesh of the environment. Neusis-NGP [44] is an extension of Neusis that leverages multi-resolution hash encodings to produce a bathymetric heightmap of the environment. To perform novel view synthesis with Neusis and Neusis-NGP, every 2D pixel that makes up a sonar image is sampled along its elevation ambiguity and then evaluated by the rendering networks to determine the intensity of each point. These points are then summed to extract the return at the specific pixel coordinate. We show that this process is time-consuming and does not produce photorealistic sonar imagery. Differentiable Space Carving (DSC) [7] proposes a faster method that models echo probabilities to carve out a mesh of the scene. This method is similar to Neusis, but avoids having to implicitly represent the scene as intensities. While this method is faster, it still requires sampling points at every pixel and then evaluating the occupancy to perform sonar novel view synthesis. Through our experiments, we show that DSC is also unable to produce high-quality novel view synthesis of imaging sonar.
Following recent developments in 3DGS, Z-Splat [32] introduces a novel method for fusing camera and sonar data, enhancing RGB image rendering by incorporating depth supervision from sonar returns. To the best of our knowledge, Z-Splat [32] is the only Gaussian splatting paper that incorporates acoustic sensor data, with demonstrations on underwater data. This is done by splatting Gaussians onto the X/Z plane and calculating a depth loss between the sonar returns and the splatted Gaussians. While this process improves the placement of Gaussians for better camera renderings, Z-Splat oversimplifies the splatting process by not accounting for any sonar-specific phenomena. More importantly, Z-Splat only splats the opacities of Gaussians, and does not consider material properties like acoustic reflectance. This leads to poor sonar renderings, which limit its application to sonar novel view synthesis. We demonstrate that SonarSplat, using range-azimuth splatting, produces more accurate novel view synthesis and can also restore sonar images to remove azimuth streaking artifacts.
3 Technical Approach
Fig. 2 shows an overview of SonarSplat, which uses sonar images and their corresponding sensor poses to optimize a scene representation. The scene is represented as a set of 3D Gaussian primitives, each parametrized with a mean, covariance, opacity, acoustic reflectivity and azimuth streaking probability. First, the Gaussians that fall within the view frustum of the sonar are rasterized into a range-azimuth image. Concurrently, we render an azimuth streak probability image using the per-Gaussian azimuth streak probabilities. Based on the relative probabilities in each range bin, we use a novel adaptive gain module to adjust the acoustic returns and capture azimuth streaking effects. Finally, we produce a sonar image with azimuth streaks that can be optimized by taking a loss with respect to the ground truth sonar image. We provide further detail of each of these components in the remainder of this section.
3.1 Gaussian Splatting Primer
This section provides background information about traditional Gaussian splatting for optical images. The standard 3D Gaussian splatting framework models the scene using a set of $N$ 3D Gaussian basis functions, $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$, each parametrized with a mean $\mu_{i}\in\mathbb{R}^{3}$, scaling $s_{i}\in\mathbb{R}^{3}$, quaternion $q_{i}\in\mathbb{R}^{4}$, opacity $o_{i}\in\mathbb{R}$, color $c_{i}\in\mathbb{R}^{3}$, encoded as spherical harmonics [13]. In this work, we augment this parametrization with acoustic-specific variables: acoustic reflectivity $\nu_{i}$ and azimuth streak probability $p_{i}$. The covariance matrix is constructed from the rotation, $q_{i}$, and scaling, $s_{i}$, of each Gaussian, each of which is converted to a rotation matrix, $R_{i}\in\mathbb{R}^{3\times 3}$, and scaling matrix, $S_{i}\in\mathbb{R}^{3\times 3}$ [13]:
$\Sigma_{i}=R_{i}S_{i}S_{i}^{T}R_{i}^{T}$ (1)
At a given camera pose $W\in SE(3)$, an image $I_{\mathcal{G}}$ can be rendered in a differentiable manner using a point-based rasterizer. The color, $C$, at a pixel, $p$, is computed by blending $\widehat{N}$ sorted Gaussians that overlap that pixel:
$C(p)=\sum_{i=1}^{\widehat{N}}c_{i}o_{i}T_{i}G_{i}$ (2)
where $T_{i}=\prod_{j=1}^{i-1}\left(1-o_{i}G_{i}\right)$ is the transmittance of the Gaussian $G_{i}$. The scene is optimized by computing the loss, $\mathcal{L}_{GS}$, between the rendered image $I_{\mathcal{G}}$ with the ground truth image $I_{\text{GT}}$, through summing $L_{1}$ loss and SSIM loss, $\mathcal{L}_{\text{SSIM}}$, as follows:
$\mathcal{L}_{GS}=\lVert I_{\mathcal{G}}(W)-I_{\text{GT}}(W)\rVert_{1}+\mathcal{L}_{\text{SSIM}}(I_{\mathcal{G}}(W),I_{\text{GT}}(W))$ (3)
3.2 Imaging Sonar Image Formation Model
To adapt the optical Gaussian splatting framework to accommodate imaging sonar, we must incorporate the acoustic image formation model. Imaging sonars emit acoustic energy from a transmitter then listen for returns on the receiver. Imaging sonar is a time-of-flight sensor that can provide the echo intensities at a given range and azimuth bin. Importantly, imaging sonar is not able to distinguish the elevation angle of a return, leading to an elevation angle ambiguity. This ambiguity makes direct use of imaging sonar difficult for 3D reconstruction and triangulation tasks [41, 11].
A sonar image is defined as $I\in\mathbb{R}^{N_{r}\times N_{a}}$ where $N_{r}$ denotes the number of range bins, and $N_{a}$ denotes the number of azimuth bins. In spherical coordinates, the sonar has a vertical field of view ($\phi_{min},\phi_{max}$), a horizontal field of view ($\theta_{min},\theta_{max}$), and a maximum range, $R_{max}$. Finally, taking $\frac{R_{max}}{N_{r}}$ gives us the range resolution, $\epsilon_{r}$, and taking $\frac{\theta_{max}-\theta{min}}{N_{a}}$ gives us the azimuthal resolution, $\epsilon_{a}$.
We leverage the sonar rendering equation introduced in [30], which calculates the intensity of return at bin $(r_{i},\theta_{j})$ in the sonar image $I$ by:
[…]
where $E_{e}$ is the intensity of emitted sound, $r$ is the range of the point, $T(r,\theta,\phi)$ is the transmittance term, and $\sigma(r,\theta,\phi)$ is the density of the point. Note that this transmittance term is equivalent to the term used in volumetric rendering such that $T$ serves as the probability that a ray travels to point $(r,\theta,\phi)$ without hitting another particle.
3.3 Range/Azimuth Splatting Model
To evaluate this integral, prior methods use sampling and quadrature [30, 31, 7]. However, motivated by the increased efficiency of splatting techniques, we explore a point-based approach adapted to the sonar image formation model. To this end, we represent the scene with a set of 3D Gaussian primitives $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$. Each of these Gaussians are parametrized with a mean $\mu_{i}\in\mathbb{R}^{3}$, scale $s_{i}\in\mathbb{R}^{3}$, orientation $q_{i}\in\mathbb{R}^{4}$, opacity $o_{i}\in\mathbb{R}$, acoustic reflectivity $\nu_{i}\in\mathbb{R}$, and azimuth streak probability $p_{i}\in\mathbb{R}$:
$G_{i}=\{\mu_{i},s_{i},q_{i},o_{i},\nu_{i},p_{i}\}$ (5)
where the covariance is computed as shown in Eq. 1.
To render sonar images, we first transform the means and covariances of the Gaussians into spherical coordinates. For a given 3D Gaussian in the sensor local frame in Cartesian coordinates, $G_{k}=\{\mu_{k}=\langle x_{k},y_{k},z_{k}\rangle,\Sigma_{k}\}$, the transformation to spherical coordinates is as follows [14]:
$\mu_{s}=\begin{bmatrix}r_{k}\\ \theta_{k}\\ \phi_{k}\end{bmatrix}=\begin{bmatrix}||\mu_{k}||\\ \arctan(y_{k},x_{k})\\ \arctan(z_{k},\sqrt{x_{k}^{2}+z_{k}^{2}})\\ \end{bmatrix}$ (6)
The covariance $\Sigma_{k}$ is transformed by using a first-order linearization of the coordinate transformation, which the Jacobian describes:
$J_{s}=\begin{bmatrix}\frac{x_{k}}{\lVert\mu_{k}\rVert}&\frac{y_{k}}{\lVert\mu_{k}\rVert}&\frac{z_{k}}{\lVert\mu_{k}\rVert}\\[4.30554pt] \frac{-y}{x^{2}+y^{2}}&\frac{x}{x^{2}+y^{2}}&0\\[4.30554pt] \frac{-x_{k}\cdot z_{k}}{||\mu_{k}||^{2}\sqrt{x^{2}+y^{2}}}&\frac{-y_{k}\cdot z_{k}}{||\mu_{k}||^{2}\sqrt{x^{2}+y^{2}}}&\frac{\sqrt{x^{2}+y^{2}}}{\lVert\mu_{k}\rVert^{2}}\end{bmatrix}$ (7)
and the covariance matrix becomes
$\Sigma_{s}=J_{s}\Sigma_{k}J_{s}^{T}$ (8)
To convert the spherical coordinate mean and covariance into range/azimuth image space, we first specify the intrinsic matrix $K$ of the sonar,
$K=\begin{bmatrix}\frac{1}{\epsilon_{r}}&0&0\\ 0&\frac{1}{\epsilon_{a}}&\frac{N_{a}}{2}\\ 0&0&1\\ \end{bmatrix}$ (9)
and then we perform the transformation
$\mu^{\prime}=K\mu_{s},\ \ \Sigma^{\prime}=J_{K}W\Sigma_{s}WJ_{K}^{T}$ (10)
where $J_{K}$ is the Jacobian of $K$, and $W$ is the sensor’s view matrix, similar to the camera splatting formulation [13].
To find the intensity of acoustic returns at a given range-azimuth bin ($r_{i},\theta_{j}$), we consider $\mathcal{G}_{i,j}\subset\mathcal{G}$, the set of Gaussians transformed to range-azimuth image space that overlap with the range-azimuth bin ($r_{i},\theta_{j}$). The rasterization equation to obtain the sonar image for SonarSplat is similar to that of [13], with a few key differences. SonarSplat’s rasterization equation is given by:
$I_{U}(r_{i},\theta_{j})=\sum_{G_{k}\in\mathcal{G}_{i,j}}\nu_{k}~{}o_{k}~{}T_{k}~{}G_{k}(r_{i},\theta_{j})$ (11)
where we introduce $\nu_{k}\in[0,1]$ as the acoustic reflectance of the Gaussian and $T_{k}$ as the transmittance from the sensor origin to the Gaussian along the range. The subscript indicates that $I_{U}$ is the unsaturated image, meaning that it does not account for the azimuth streaking that causes saturation across the image. We observe that the acoustic reflectance of a set of points depends on material properties and viewing angle, so $\nu_{k}$ uses spherical harmonics to encode view-dependent effects.
Figure 3: Azimuth streaking is characterized by black streaks (referred to by blue arrows) spanning the azimuth axis of a sonar image. (a) Sonar image of a metal frame structure that experiences azimuth streaking. (b) Sonar image of a concrete piling that has multiple azimuth streaks. Image brightness was enhanced for illustrative purposes. Images are polar (range/azimuth).
3.4 Modeling Azimuth Streaking
Azimuth streaking, as shown in Fig. 3, is a phenomena that is frequently observed in sonar images. It is a form of saturation that occurs when the sonar receiver receives strong returns from incident angles close to parallel at a specific range bin $r_{i}$ [29]. The azimuth streaking model found in [29] proposes to modify all the returns in a range interval if a certain percentage of returns exceeds a given threshold $\zeta$. The image with added azimuth streaks is given by:
$\hat{I}(r_{i},\cdot)=1-(1-I(r_{i},\cdot))^{2}$ (12)
where $\hat{I}(r_{i},\cdot)$ is the streaked image, and $I(r_{i},\cdot)$ is the vector of intensities at range interval $r_{i}$.
In our model, we assign each Gaussian a probability that it will cause an azimuth streak. During the splatting process, we accumulate this probability as a function of the Gaussian’s opacity, mean, and covariance. The benefit of defining this probability per-Gaussian is that it can be optimized to produce azimuth streaks that are multi-view consistent.
Let $p_{k}\in[0,1]$ be the probability that Gaussian $k$ will contribute to an azimuth streaking phenomena. We can then splat these per-Gaussian probabilities into a range-azimuth bin to obtain an azimuth streak probability image $P_{a}$:
$P_{a}(r_{i},\theta_{j})=\sum_{G_{k}\in\mathcal{G}_{i,j}}p_{k}~{}o_{k}~{}T_{k}~{}G_{k}(r_{i},\theta_{j})$ (13)
$M_{a}(r_{i})=\sum_{\theta=0}^{N_{a}}P_{a}(r_{i},\theta)$ (14)
where $M_{a}(r_{i})$ represents the probability of an azimuth streak occurring at range interval $r_{i}$. The azimuth streak probabilities across the range bins are then used to compute the final image using an adaptive gain mechanism.
3.4.1 Adaptive Gain Module
We introduce a novel adaptive gain term, $A$, to transform the azimuth streak probability image into receiver gains. We begin designing this gain term with a few observations. First, when no Gaussian in a range bin $r_{i}$ has a high probability of azimuth streaking, the gain should be unity. Secondly, suppose a single Gaussian $G_{k}$ exhibits a high probability of azimuth streaking across the range bin $r_{i}$. In that case, we wish to adaptively suppress the other Gaussians in $r_{i}$ and assign a high gain to $G_{k}$. Finally, if multiple Gaussians in $r_{i}$ have high probabilities of contributing to an azimuth streak, higher gain should be assigned to those Gaussians. To address these observations, the adaptive gain term is applied to each range-azimuth bin $(r_{i},\theta_{j})$:
$A(r_{i},\theta_{j})=P_{a}(r_{i},\theta_{j})M_{a}(r_{i})\frac{e^{\gamma\cdot P_{a}(r_{i},\theta_{j})}+1}{e^{\gamma}+1}+(1-M_{a}(r_{i}))$
where $\gamma$ is a scaling factor that dictates the steepness of the adaptive gain. This behavior aligns with the insight presented in [29]: if a certain percentage of returns exceeds a threshold, an operation is applied to all the values in a range interval. However, our proposed model differs because we do not restrict the gain to a fixed function (quadratic) but rather offer a family of curves for our splatting model to explore during optimization. Further discussion on the adaptive gain term can be found in the supplementary material.
The final image with the azimuth streaks is then computed by applying the gain $A(r_{i},\theta_{j})$ to its corresponding bin on $I_{U}(r_{i},\theta_{j})$:
$\hat{I}(r_{i},\theta_{j})=A(r_{i},\theta_{j})\cdot I_{U}(r_{i},\theta_{j})$ (16)
resulting in the final rendered sonar image $\hat{I}(r_{i},\theta_{j})$.
Figure 4: We present selected images from our diverse set of 8 sequences used to evaluate of our method. We show RGB images of the environment for visualization, and captured range/azimuth sonar images. Our evaluation datasets focus on smaller objects and larger-scale structures such as the test-tank basin. Sonar images are polar (range/azimuth).
3.5 Model Optimization
We optimize for desired parameters using SGD and by taking losses ($\mathcal{L}_{l1},\mathcal{L}_{ssim}$) between $\hat{I}$ and the ground truth image $I_{\text{GT}}$[13]. To properly optimize for the azimuth streaking probabilities $p_{k}$ of each Gaussian, we first identify where azimuth streaks occur by calculating the average intensity of the range interval $r_{i}$. Then, for the first $N_{s}$ iterations, we train only on pixels in range bins that exceed a certain average intensity threshold $\tau$. Importantly, $p_{k}$ is not optimized during this interval. Then, after $N_{s}$ iterations, we train only on pixels less than $\tau$. We do not optimize $\mu_{k},\Sigma_{k},o_{k}$, or $r_{k}$ during this interval. This way, we allow the Gaussians to initialize and fit the images before isolating and optimizing the azimuth streaking parameters.
We also encourage each Gaussian to have either high or low opacity via an opacity loss $\mathcal{L}_{o}$. This prevents a set of medium-opacity Gaussians from producing the same pixel intensity. Rather, the model is encouraged to use the reflectance parameter $r_{k}$ to represent low-intensity returns. We use the negative log-likelihood of two Laplacian distributions as in [15].
$\mathcal{L}_{o}(x)=|x||1-x|$ (17)
The final loss becomes:
$\mathcal{L}=\lambda_{l1}\mathcal{L}_{l1}+\lambda_{ssim}\mathcal{L}_{ssim}+\lambda_{o}\mathcal{L}_{o}$ (18)
where $\lambda_{l1},\lambda_{ssim},\text{and}~{}\lambda_{o}$ are the weights for the L1, SSIM, and opacity losses respectively.
4 Results & Experiments
4.1 Datasets
We note that real, public datasets of posed sonar images suitable for novel view synthesis experiments do not exist. Rather, we collect sequences using our robotic platform based on a BlueRobotics BlueROV2 vehicle shown in Fig. 1. Our robotic platform is equipped with a BluePrint Subsea Oculus M750-d imaging sonar capable of capturing data across various ranges with a 130 horizontal field of view and a 20 vertical field of view. We use this platform to collect 8 trajectories inspecting a range of objects and in both a test tank and river environment. Fig. 4 shows selected sequences for visualization. For detailed information about sonar parameters, sensor pose estimation, our platform, and dataset details, please refer to the supplemental material.
4.2 Baselines
For baselines, we compare to Neusis [30], Neusis-NGP [44], DSC [7], and ZSplat*. Note that ZSplat [32] was originally designed for camera-sonar fusion. Since we are focused on sonar-only rendering, we compare to a modified version, ZSplat*, trained with the sonar loss.
Figure 5: We present qualitative novel view synthesis results from selected datasets and baselines. Note that SonarSplat produces superior sonar image synthesis compared to baselines and is better able to capture finer details in the environment. SonarSplat is also able to capture azimuth streaking phenomena, which no other baseline is able to do. We increase the brightness for all outputs by 25% for easier viewing. Note that Z-Splat* indicates Z-Splat trained only with the sonar loss. Images are cartesian.
De-Streaking Performance (ICV) ($\uparrow$)
Sequence Captured Image SonarSplat NVS De-Streaked Image
basin_empty 2.74 9.78 19.44
infra_360 2.89 34.30 44.18
piling_1 2.78 11.47 19.51
basin_infra 2.77 4.41 4.84
rock_360 2.86 25.00 44.47
piling_2 2.79 7.61 10.32
pole_river 3.11 2.69 2.78
rock_river 3.15 5.15 6.78
Average 2.89 12.55 19.04
Table 1: Quantitative results for azimuth streak removal. We report the inverse coefficient of variation (ICV) [37]. Higher is better ($\uparrow$), and best is in bold.
4.3 Novel View Synthesis
We train our method on an NVIDIA RTX 3090 GPU with 24 GB of vRAM. For detailed implementation details, please refer to the supplemental material. We report validation PSNR, SSIM, and LPIPS following [13, 21]. For our validation set, we take every 8th image from the dataset as introduced in [21]. We report SonarSplat’s performance compared to baselines in Tab. 2. We find that SonarSplat consistently synthesizes more realistic sonar images, quantified by its higher PSNR, SSIM, and lower LPIPS values. We also present qualitative results from selected sequences in Fig. 5. Note that these qualitative results are in cartesian space for interpretability. For extensive polar image results, please refer to the supplemental material. Finally, we explore the feasibility of SonarSplat for data synthesis applications by reporting the rendering speed in frames per second (FPS) of each method in Tab. 3. We calculate the FPS of each method on an NVIDIA A6000 GPU with 50 GB of vRAM. SonarSplat and ZSplat both leverage Gaussian splatting, which offers significantly faster rendering speeds compared to neural rendering methods. ZSplat is optimized for high rendering speed, which will be a focus of future work for SonarSplat.
Sonar Image Synthesis
Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
Neusis [30] 14.43 0.03 0.66
Neusis-NGP [44] 19.43 0.33 0.55
DSC [7] 14.87 0.10 0.63
Z-Splat* [32] 8.57 0.03 0.68
SonarSplat (ours) 21.89 0.41 0.62
Table 2: Novel view synthesis performance on real world data. Note that ZSplat* is ZSplat with only the sonar loss activated for training. Best is shown in bold.
Figure 6: Qualitative results for azimuth streak removal capabilities of SonarSplat. Note that other baselines are unable to remove azimuth streaks from synthesized images since they do not learn the streaking parameters. Locations of azimuth streaks in the captured images are shown in blue arrows for illustrative purposes. SonarSplat is able to effectively restore sonar imagery that exhibits azimuth streaking. Images are cartesian.
Rendering Speed ($\uparrow$)
Method FPS
Neusis [30] 0.03
Neusis-NGP [44] 0.01
DSC [7] 0.03
Z-Splat* [32] 155.32
SonarSplat (ours) 9.46
Table 3: Rendering speed in frames per second (FPS) averaged across all validation images across all sequences. All results are reported on an NVIDIA RTX A6000 GPU. Note that ZSplat* is ZSplat with only sonar rendering. Best is shown in bold.
Figure 7: Qualitative results from 3D reconstruction of SonarSplat compared to baselines.
4.4 Azimuth Streak Removal
Since SonarSplat learns probabilities of azimuth streaking for each Gaussian, we can undo the adaptive gain to render images without azimuth streaks present and recover suppressed returns. This vision task is most similar to image de-striping from the remote sensing literature. We present quantitative results in Tab. 1 using the inverse coefficient of variation (ICV) metric, which is a common no-reference metric used in de-striping works [36, 33, 26, 37]. We present qualitative results of de-streaked images in Fig. 6.
4.5 3D Reconstruction
We show that SonarSplat learns the geometric properties of the scene by converting the optimized splat into a pointcloud for visualization. We compare with methods focusing on 3D reconstruction: Neusis and DSC [30, 7]. Qualitative results are shown in Fig. 7 and quantitative results are shown in Tab. 4. We use popular metrics from scene reconstruction literature and choose $l1$ Chamfer Distance, F-Score, and voxelized IoU to evaluate performance [1]. For Neusis and DSC, we produce meshes using marching cubes. For SonarSplat, we choose to sample points from the 3D Gaussians to produce a pointcloud. We sample 100,000 points from both the mesh and Gaussian splat and compute evaluation metrics against a ground truth photogrammetry reconstruction of the rock platform. We note that neither Neusis nor DSC are able to reconstruct the object whereas SonarSplat is able to produce a coarse representation.
Reconstruction Metrics
Method CD-$l_{1}$ (m) $\downarrow$ F-Score $\uparrow$ IoU $\uparrow$
Neusis [30] 0.202 0.4834 0.0129
DSC [7] 0.187 0.4613 0.0164
SonarSplat (ours) 0.084 0.1170 0.0191
Table 4: We quantitatively evaluate the 3D reconstruction of selected methods on the rock_360 sequence. Best is in bold.
5 Conclusion & Future Work
We propose SonarSplat, the first sonar-only Gaussian splatting framework for novel view synthesis for imaging sonar in underwater applications. First, we adapt the sonar rendering equation for efficient range-azimuth splatting by evaluating 3D Gaussians. We model azimuth streaking within the Gaussian splatting framework, allowing us to learn per-Gaussian azimuth streaking probabilities and produce high-quality de-streaked images. We perform experiments on real-world data in test tank and river environments. In the novel view synthesis task, SonarSplat outperforms baselines by +2.5 dB PSNR and demonstrates superior qualitative results. We also demonstrate that SonarSplat learns scene geometry, demonstrated by both quantitative and qualitative results. Future work will focus on leveraging SonarSplat for sonar data synthesis by randomizing scene parameters including acoustic reflectance and azimuth streaking probability. This can yield more diverse datasets for training sonar-based scene understanding algorithms. We will also work to improve SonarSplat’s ability to perform dense 3D reconstruction by better modeling additional acoustic phenomena like speckle noise, Lambertian scattering, and multi-path reflections.
6 Acknowledgment
This work is supported by the National Science Foundation under Award No. 2337774.

### v2

@@
 Our method represents the scene as a set of 3D Gaussians with acoustic reflectance and saturation properties.
-We develop a novel method to efficiently rasterize learned Gaussians to produce a range/azimuth image that is faithful to the acoustic image formation model of imaging sonar.
+We develop a novel method to efficiently rasterize Gaussians to produce a range/azimuth image that is faithful to the acoustic image formation model of imaging sonar.
 In particular, we develop a novel approach to model azimuth streaking in a Gaussian splatting framework.
 We evaluate SonarSplat using real-world datasets of sonar images collected from an underwater robotic platform in a controlled test tank and in a real-world river environment.
-Compared to the state-of-the-art, SonarSplat offers improved image synthesis capabilities (+2.5 dB PSNR).
-We also demonstrate that SonarSplat can be leveraged for azimuth streak removal and 3D scene reconstruction.
-1 Introduction
-Figure 1: We present SonarSplat, a method that uses posed sonar images to produce a novel 3D Gaussian Splat representation that enables novel view synthesis of realistic sonar images and coarse 3D reconstruction.
-With SonarSplat, the scene is represented with a set of 3D Gaussians that model the reflectivity and azimuth streaking probabilities of objects in the rendering process.
-This novel formulation enables additional features of SonarSplat, such as azimuth streak removal, enabling the recovery of clean sonar images from saturated ones.
+Compared to the state-of-the-art, SonarSplat offers improved image synthesis capabilities (+3.2 dB PSNR) and more accurate 3D reconstruction (52% lower Chamfer Distance).
+We also demonstrate that SonarSplat can be leveraged for azimuth streak removal.
+I Introduction
 Acoustic sensors, such as imaging sonar, are commonly used for infrastructure inspection, large-area mapping, and target detection in underwater environments [12, 39, 24].
@@
 Although sonar exhibits many desirable qualities such as longer range, invariance to lighting conditions, and the ability to discern material properties, there exist acoustic phenomena, including elevation ambiguity, azimuth streaking, and multi-path reflections, which make sonar interpretation difficult for operators and computer vision algorithms alike.
-Furthermore, the severe lack of large, public datasets for sonar slows progress in traditional computer vision tasks like object detection, segmentation, and reconstruction.
-Recently, neural radiance fields (NeRFs) [22] have demonstrated the potential for high-fidelity data synthesis and denoising of optical imagery [27, 6].
+Figure 1: We present SonarSplat, a novel 3D Gaussian splatting framework that enables novel view synthesis of realistic sonar images, image de-streaking, and 3D reconstruction.
+We model the acoustic reflectivity and azimuth streak probabilities for all Gaussians, yielding a 3D scene representation that can be queried for a variety of downstream tasks.
+Radiance field methods offer a promising approach to addressing these challenges.
+Neural radiance fields (NeRFs) have demonstrated the potential for high-fidelity data synthesis, denoising, and 3D reconstruction using optical imagery [26, 45, 30, 5].
 The benefits of neural rendering have been recognized by the underwater perception community, with prior work exploring neural rendering for 3D object reconstruction using underwater cameras [18, 41] and sonar data [33, 34, 50, 6, 49].
-While these innovations are promising, training and processing NeRFs is extremely costly and time-consuming, making deploying them in real-time or on resource-constrained devices difficult.
+While these innovations are promising, training and processing NeRFs is costly and time-consuming, making deploying them in real-time or on resource-constrained devices difficult.
 More recently, 3D Gaussian splatting (3DGS) [14] was developed as a faster alternative to NeRF.
-3DGS utilizes sparse points in the form of 3D Gaussians to represent the scene.
-This allows 3DGS to preserve properties of continuous volumetric radiance fields while lowering computational costs.
-Gaussian splatting has been leveraged for underwater imagery [45, 16, 23].
+Gaussian splatting has been leveraged for underwater applications as well [51, 19, 27].
 Most relevant to our proposed work, ZSplat proposes a Gaussian splatting framework for RGB-sonar fusion [35].
-However, ZSplat leverages the fusion of sonar data to improve the rendering of RGB images and does not focus on enabling high-fidelity data synthesis for sonar imagery or provide evaluation for the quality of rendered sonar images.
-Thus, there is a clear gap in the literature for a framework capable of efficient and effective sonar image synthesis.
+However, ZSplat leverages the fusion of sonar data to improve the rendering of RGB images and does not focus on enabling high-fidelity data synthesis for sonar imagery or provide evaluation for the quality of rendered sonar images or 3D reconstruction.
+Thus, there is a clear gap in the literature for a framework capable of efficient and effective sonar image synthesis and 3D reconstruction using Gaussian splatting.
 In this paper, we propose SonarSplat (Fig.
-1), a novel Gaussian splatting framework for imaging sonar that enables efficient and high quality novel view synthesis for imaging sonar data.
-We focus on modeling characteristic acoustic noise properties, specifically, azimuth streaking, or receiver saturation, which can make image interpretation and 3D reconstruction difficult.
-By modeling realistic noise effects, we demonstrate that we can leverage SonarSplat to perform efficient, high quality sonar image synthesis and noise removal.
-Lastly, we also demonstrate SonarSplat’s ability to learn geometry of the scene, overcoming the elevation ambiguity inherent in imaging sonar image formation.
+1), a novel Gaussian splatting framework for imaging sonar that enables efficient and high quality novel view synthesis, sonar image de-streaking, and 3D reconstruction for underwater robotic applications.
+To the best of our knowledge, SonarSplat is the first sonar-only method that leverages Gaussian splatting to perform these tasks.
 Our main contributions are summarized as follows:
@@
   • We develop a novel method for learning the probability of azimuth streaking in a differentiable manner, which allows for removal of unwanted azimuth streaking artifacts from rendered images.
+  • We introduce a novel densification strategy called the Elevation Sampling Densification Strategy (ESDS) that places Gaussians on elevation arcs and yields better novel view synthesis and 3D reconstruction.
   • We perform quantitative and qualitative evaluation of our method on real imaging sonar data collected from a robotic platform with a variety of sonar ranges and environments.
-Code and data will be made publicly available on a project website after a double-blind review.
-2 Related Work
+Code and data will be made publicly available on a project website after review.
+II Related Work
 Figure 2: Overview of SonarSplat.
@@
 Additionally, we optimize and splat per-Gaussian azimuth streaking probabilities $p_{k}$.
-All the probabilities in range interval $r_{i}$ are considered in our novel Adaptive Gain Module, which adjusts the receiver gain applied to $I_{u}$.
+All of the probabilities in a range bin $r_{i}$ are considered in our novel Adaptive Gain Module, which adjusts the receiver gain applied to $I_{u}$.
 Finally, we produce $\hat{I}$ by multiplying the gain $A$ by $I_{u}$.
 All parameters are optimized using gradient descent by taking losses with respect to the sonar image pixel values.
-All sonar images shown are polar (range-azimuth) coordinates.
-2.1 Radiance Field Methods
-NeRFs represent scenes implicitly using neural networks and have been explored for a variety of applications, including novel view synthesis, inverse rendering, data generation, and image denoising [21, 8, 27, 6].
-Later works improved NeRFs through improved sampling of the volumetric representation and improvements in rendering speed with alternative encoding methods [24, 2].
-Despite their vast success as a learned scene representation, a key struggle for the practical application of NeRFs is the long training time and the slow rendering speeds, primarily due to the computationally expensive volumetric rendering process [21].
-A recent development in radiance field methods is Gaussian splatting, which uses rasterization instead of volumetric rendering to optimize and produce a scene representation [13].
-The resulting renders from 3DGS are both high in quality and reach unprecedented rendering speeds, enabling fast rendering for various real-time applications [19, 18].
+All sonar images shown are polar (range/azimuth) coordinates.
+II-A Radiance Field Methods
+NeRFs represent scenes implicitly using neural networks and have been explored for a variety of robotic applications, including novel view synthesis, data generation, Simultaneous Localization and Mapping (SLAM), manipulation, image dehazing, and trajectory optimization [26, 2, 11, 37, 43, 5, 7].
+Despite their success as a scene representation, a key challenge for the practical application of NeRFs is the long training time and the slow rendering speeds, primarily due to the computationally expensive volumetric rendering process [26].
+A recent development in radiance field methods is Gaussian splatting, which uses more efficient rasterization instead of volumetric rendering to optimize and produce a scene representation, enabling usage in a variety of real-time applications [14, 23, 22, 25].
 Despite their impressive performance and rapid adoption, radiance field methods were developed with optical sensors in mind, limiting their direct application to non-visual sensors.
 Recent developments have integrated the physics-based image formation models of various non-visual sensors, including radar, LIDAR, sonar, thermal cameras, and event cameras into NeRF and Gaussian splatting frameworks for various vision tasks [3, 4, 9, 33, 20].
-In this work, we propose a novel Gaussian splatting framework for imaging sonar for underwater applications.
-2.2 Perception Challenges for Imaging Sonar
+We are inspired by the faster rendering speeds and high fidelity of Gaussian splatting methods and propose a novel Gaussian splatting framework for imaging sonar.
+II-B Perception Challenges for Imaging Sonar
 Imaging sonar is a popular underwater sensor for perception tasks like object detection, segmentation, reconstruction, and inspection [33, 8, 12, 44, 28].
-The key challenges with working with imaging sonars for these tasks are properties inherent with the acoustic image formation model [29].
-The sensor is susceptible to multi-path effects, highly nonlinear noise patterns, and phenomena like azimuth streaking [28].
+However, the sensor is susceptible to multi-path effects, nonlinear noise patterns, and phenomena like azimuth streaking [31].
 Azimuth streaking, in particular, can make sonar image understanding difficult by suppressing returns from certain azimuth angles, causing black bars in the sonar image [31].
-In addition, the sensor inherently loses information from 3D into 2D when capturing the acoustic returns, causing an effect known as elevation ambiguity [40].
-This ambiguity has made the successful application of imaging sonar to 3D vision tasks difficult [25, 11, 42].
-Yet, imaging sonars still remain an extremely valuable sensor for underwater imaging, especially in long range imaging and deep water applications where optical sensors may fail [39].
-Given its importance to underwater perception, prior work has proposed sonar simulators to validate perception algorithms without the cost, time commitment, and difficulty of real-world data collection [29, 5].
-Although these works model complex interactions with the environment, such as multi-path reflections and azimuth streaking, there still exists a sim-to-real gap inherent with simulators.
-Furthermore, simulators also require the manual tuning of acoustic and scene parameters to match the distribution of data desired by the end-user.
-Rather, our proposed work proposes a Gaussian splatting framework that can learn from real data to produce realistic sonar data from novel poses, which may be useful for augmenting datasets and addressing the data scarcity problem.
-2.3 Novel View Synthesis for Imaging Sonar
-Prior work has incorporated the sonar image formation model into signed distance field (SDF) and NeRF formulations [30, 7].
-Neusis [30] presents a NeRF framework for dense 3D reconstruction of objects using an imaging sonar, specifically, a forward-looking sonar (FLS).
-Neusis proposes a differentiable volumetric renderer to model the propagation of acoustic waves to synthesize imaging sonar measurements.
-By using this implicit representation and multiple sensor measurements, Neusis attempts to resolve the elevation ambiguity present in the image data to construct a 3D mesh of the environment.
-Neusis-NGP [44] is an extension of Neusis that leverages multi-resolution hash encodings to produce a bathymetric heightmap of the environment.
-To perform novel view synthesis with Neusis and Neusis-NGP, every 2D pixel that makes up a sonar image is sampled along its elevation ambiguity and then evaluated by the rendering networks to determine the intensity of each point.
-These points are then summed to extract the return at the specific pixel coordinate.
-We show that this process is time-consuming and does not produce photorealistic sonar imagery.
-Differentiable Space Carving (DSC) [7] proposes a faster method that models echo probabilities to carve out a mesh of the scene.
-This method is similar to Neusis, but avoids having to implicitly represent the scene as intensities.
-While this method is faster, it still requires sampling points at every pixel and then evaluating the occupancy to perform sonar novel view synthesis.
-Through our experiments, we show that DSC is also unable to produce high-quality novel view synthesis of imaging sonar.
-Following recent developments in 3DGS, Z-Splat [32] introduces a novel method for fusing camera and sonar data, enhancing RGB image rendering by incorporating depth supervision from sonar returns.
-To the best of our knowledge, Z-Splat [32] is the only Gaussian splatting paper that incorporates acoustic sensor data, with demonstrations on underwater data.
-This is done by splatting Gaussians onto the X/Z plane and calculating a depth loss between the sonar returns and the splatted Gaussians.
-While this process improves the placement of Gaussians for better camera renderings, Z-Splat oversimplifies the splatting process by not accounting for any sonar-specific phenomena.
-More importantly, Z-Splat only splats the opacities of Gaussians, and does not consider material properties like acoustic reflectance.
-This leads to poor sonar renderings, which limit its application to sonar novel view synthesis.
-We demonstrate that SonarSplat, using range-azimuth splatting, produces more accurate novel view synthesis and can also restore sonar images to remove azimuth streaking artifacts.
-3 Technical Approach
-Fig.
-2 shows an overview of SonarSplat, which uses sonar images and their corresponding sensor poses to optimize a scene representation.
+In addition, the sensor inherently loses information from 3D into 2D when capturing the acoustic returns, causing an effect known as elevation angle ambiguity [46].
+This ambiguity has made the successful application of imaging sonar to 3D vision tasks challenging [28, 10, 48].
+II-C Novel View Synthesis for Imaging Sonar
+Prior work has incorporated the sonar image formation model into signed distance field (SDF) and NeRF formulations [33, 6, 50].
+Neusis presents a volumetric rendering framework for dense 3D reconstruction of objects using an imaging sonar [33].
+Neusis-NGP is an extension of Neusis that leverages multi-resolution hash encodings to produce a bathymetric heightmap of the environment [50].
+Differentiable Space Carving (DSC) proposes a faster method that models echo probabilities to carve out a mesh of the scene [6].
+Through our experiments, we show that none of these method are able to perform realistic looking novel view synthesis for imaging sonar, limiting their application to data generation and augmentation.
+Furthermore, querying these methods is time consuming and limits their application to real-time methods.
+To the best of our knowledge, ZSplat is the only Gaussian splatting paper that incorporates imaging sonar data, with demonstrations on underwater data.
+However, ZSplat oversimplifies the splatting process by not accounting for any sonar-specific phenomena.
+More importantly, ZSplat only splats the opacities of Gaussians and does not consider material properties like acoustic reflectance.
+This leads to poor sonar renderings, which limits its application to sonar novel view synthesis.
+We demonstrate that SonarSplat, which more accurately models sonar image formation, produces more realistic novel view synthesis, restores sonar images to remove azimuth streaking artifacts, and produces 3D reconstructions under a unified framework.
+III Technical Approach
+Figure 2 shows an overview of SonarSplat, which uses sonar images and their corresponding sensor poses to optimize a scene representation.
 The scene is represented as a set of 3D Gaussian primitives, each parametrized with a mean, covariance, opacity, acoustic reflectivity and azimuth streaking probability.
-First, the Gaussians that fall within the view frustum of the sonar are rasterized into a range-azimuth image.
+First, the Gaussians that fall within the view frustum of the sonar are rasterized into a range/azimuth image.
 Concurrently, we render an azimuth streak probability image using the per-Gaussian azimuth streak probabilities.
@@
 We provide further detail of each of these components in the remainder of this section.
-3.1 Gaussian Splatting Primer
-This section provides background information about traditional Gaussian splatting for optical images.
-The standard 3D Gaussian splatting framework models the scene using a set of $N$ 3D Gaussian basis functions, $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$, each parametrized with a mean $\mu_{i}\in\mathbb{R}^{3}$, scaling $s_{i}\in\mathbb{R}^{3}$, quaternion $q_{i}\in\mathbb{R}^{4}$, opacity $o_{i}\in\mathbb{R}$, color $c_{i}\in\mathbb{R}^{3}$, encoded as spherical harmonics [13].
-In this work, we augment this parametrization with acoustic-specific variables: acoustic reflectivity $\nu_{i}$ and azimuth streak probability $p_{i}$.
-The covariance matrix is constructed from the rotation, $q_{i}$, and scaling, $s_{i}$, of each Gaussian, each of which is converted to a rotation matrix, $R_{i}\in\mathbb{R}^{3\times 3}$, and scaling matrix, $S_{i}\in\mathbb{R}^{3\times 3}$ [13]:
-$\Sigma_{i}=R_{i}S_{i}S_{i}^{T}R_{i}^{T}$ (1)
-At a given camera pose $W\in SE(3)$, an image $I_{\mathcal{G}}$ can be rendered in a differentiable manner using a point-based rasterizer.
-The color, $C$, at a pixel, $p$, is computed by blending $\widehat{N}$ sorted Gaussians that overlap that pixel:
-$C(p)=\sum_{i=1}^{\widehat{N}}c_{i}o_{i}T_{i}G_{i}$ (2)
-where $T_{i}=\prod_{j=1}^{i-1}\left(1-o_{i}G_{i}\right)$ is the transmittance of the Gaussian $G_{i}$.
-The scene is optimized by computing the loss, $\mathcal{L}_{GS}$, between the rendered image $I_{\mathcal{G}}$ with the ground truth image $I_{\text{GT}}$, through summing $L_{1}$ loss and SSIM loss, $\mathcal{L}_{\text{SSIM}}$, as follows:
-$\mathcal{L}_{GS}=\lVert I_{\mathcal{G}}(W)-I_{\text{GT}}(W)\rVert_{1}+\mathcal{L}_{\text{SSIM}}(I_{\mathcal{G}}(W),I_{\text{GT}}(W))$ (3)
-3.2 Imaging Sonar Image Formation Model
+III-A Imaging Sonar Formation Model
 To adapt the optical Gaussian splatting framework to accommodate imaging sonar, we must incorporate the acoustic image formation model.
-Imaging sonars emit acoustic energy from a transmitter then listen for returns on the receiver.
 Imaging sonar is a time-of-flight sensor that can provide the echo intensities at a given range and azimuth bin.
@@
 This ambiguity makes direct use of imaging sonar difficult for 3D reconstruction and triangulation tasks [47, 10].
-A sonar image is defined as $I\in\mathbb{R}^{N_{r}\times N_{a}}$ where $N_{r}$ denotes the number of range bins, and $N_{a}$ denotes the number of azimuth bins.
+A sonar image is defined as $I\in\mathbb{R}^{N_{a}\times N_{r}}$ where $N_{a}$ denotes the number of azimuth bins and $N_{r}$ denotes the number of range bins.
 In spherical coordinates, the sonar has a vertical field of view ($\phi_{min},\phi_{max}$), a horizontal field of view ($\theta_{min},\theta_{max}$), and a maximum range, $R_{max}$.
@@
 where $E_{e}$ is the intensity of emitted sound, $r$ is the range of the point, $T(r,\theta,\phi)$ is the transmittance term, and $\sigma(r,\theta,\phi)$ is the density of the point.
-Note that this transmittance term is equivalent to the term used in volumetric rendering such that $T$ serves as the probability that a ray travels to point $(r,\theta,\phi)$ without hitting another particle.
-3.3 Range/Azimuth Splatting Model
+Note that this transmittance term is equivalent to the term used in volumetric rendering, serving as the probability that a ray travels to point $(r,\theta,\phi)$ without hitting another particle.
+III-B Range/Azimuth Splatting Model
 To evaluate this integral, prior methods use sampling and quadrature [33, 34, 6].
 However, motivated by the increased efficiency of splatting techniques, we explore a point-based approach adapted to the sonar image formation model.
-To this end, we represent the scene with a set of 3D Gaussian primitives $\mathcal{G}=\{G_{i}\}_{i=1}^{N}$.
-Each of these Gaussians are parametrized with a mean $\mu_{i}\in\mathbb{R}^{3}$, scale $s_{i}\in\mathbb{R}^{3}$, orientation $q_{i}\in\mathbb{R}^{4}$, opacity $o_{i}\in\mathbb{R}$, acoustic reflectivity $\nu_{i}\in\mathbb{R}$, and azimuth streak probability $p_{i}\in\mathbb{R}$:
-$G_{i}=\{\mu_{i},s_{i},q_{i},o_{i},\nu_{i},p_{i}\}$ (5)
-where the covariance is computed as shown in Eq. 1.
+To this end, we represent the scene with a set of 3D Gaussian primitives $\mathcal{G}=\{G_{k}\}_{k=1}^{N}$.
+We augment this parametrization with acoustic-specific variables: acoustic reflectivity $\nu_{k}$ and azimuth streak probability $p_{k}$.
+Formally, these Gaussians are parametrized with a mean $\mu_{k}\in\mathbb{R}^{3}$, scale $s_{k}\in\mathbb{R}^{3}$, orientation $q_{k}\in\mathbb{H}$, opacity $o_{k}\in\mathbb{R}^{+}$, acoustic reflectivity $\nu_{k}\in\mathbb{R}^{+}$, and azimuth streak probability $p_{k}\in\mathbb{R}^{+}$:
+$G_{k}=\{\mu_{k},s_{k},q_{k},o_{k},\nu_{k},p_{k}\}$ (2)
+where the covariance $\Sigma_{k}$ is computed as in [14].
 To render sonar images, we first transform the means and covariances of the Gaussians into spherical coordinates.
-For a given 3D Gaussian in the sensor local frame in Cartesian coordinates, $G_{k}=\{\mu_{k}=\langle x_{k},y_{k},z_{k}\rangle,\Sigma_{k}\}$, the transformation to spherical coordinates is as follows [14]:
-$\mu_{s}=\begin{bmatrix}r_{k}\\ \theta_{k}\\ \phi_{k}\end{bmatrix}=\begin{bmatrix}||\mu_{k}||\\ \arctan(y_{k},x_{k})\\ \arctan(z_{k},\sqrt{x_{k}^{2}+z_{k}^{2}})\\ \end{bmatrix}$ (6)
-The covariance $\Sigma_{k}$ is transformed by using a first-order linearization of the coordinate transformation, which the Jacobian describes:
+Let $\mu_{k}=\langle x_{k},y_{k},z_{k}\rangle~{}\text{and}~{}\Sigma_{k}$ be the mean and covariance in the sensor’s frame, respectively.
+The transformation to spherical coordinates is as follows [15]:
+$\mu_{s}=\begin{bmatrix}r_{k}\\ \theta_{k}\\ \phi_{k}\end{bmatrix}=\begin{bmatrix}||\mu_{k}||\\ \arctan(y_{k},x_{k})\\ \arctan(z_{k},\sqrt{x_{k}^{2}+y_{k}^{2}})\\ \end{bmatrix}$ (3)
+The covariance matrix $\Sigma_{k}$ is transformed by using a first-order linearization of the coordinate transformation, which the Jacobian describes:
 $J_{s}=\begin{bmatrix}\frac{x_{k}}{\lVert\mu_{k}\rVert}&\frac{y_{k}}{\lVert\mu_{k}\rVert}&\frac{z_{k}}{\lVert\mu_{k}\rVert}\\[4.30554pt] \frac{-y}{x^{2}+y^{2}}&\frac{x}{x^{2}+y^{2}}&0\\[4.30554pt] \frac{-x_{k}\cdot z_{k}}{||\mu_{k}||^{2}\sqrt{x^{2}+y^{2}}}&\frac{-y_{k}\cdot z_{k}}{||\mu_{k}||^{2}\sqrt{x^{2}+y^{2}}}&\frac{\sqrt{x^{2}+y^{2}}}{\lVert\mu_{k}\rVert^{2}}\end{bmatrix}$ (4)
@@
 and then we perform the transformation
-$\mu^{\prime}=K\mu_{s},\ \ \Sigma^{\prime}=J_{K}W\Sigma_{s}WJ_{K}^{T}$ (10)
-where $J_{K}$ is the Jacobian of $K$, and $W$ is the sensor’s view matrix, similar to the camera splatting formulation [13].
-To find the intensity of acoustic returns at a given range-azimuth bin ($r_{i},\theta_{j}$), we consider $\mathcal{G}_{i,j}\subset\mathcal{G}$, the set of Gaussians transformed to range-azimuth image space that overlap with the range-azimuth bin ($r_{i},\theta_{j}$).
+$\mu_{k}^{\prime}=K\mu_{s},\ \ \Sigma_{k}^{\prime}=J_{K}W\Sigma_{s}WJ_{K}^{T}$ (7)
+where $\mu_{k}^{\prime}$ and $\Sigma_{k}^{\prime}$ are the mean and covariance transformed to range/azimuth space, $J_{K}$ is the Jacobian of $K$, and $W$ is the sensor’s view matrix, similar to the camera splatting formulation [14].
+To find the intensity of acoustic returns at a given range/azimuth bin ($r_{i},\theta_{j}$), we consider $\mathcal{G}_{i,j}\subset\mathcal{G}$, the set of Gaussians transformed to range/azimuth image space that overlap with the range/azimuth bin ($r_{i},\theta_{j}$).
 The rasterization equation to obtain the sonar image for SonarSplat is similar to that of [14], with a few key differences.
 SonarSplat’s rasterization equation is given by:
-$I_{U}(r_{i},\theta_{j})=\sum_{G_{k}\in\mathcal{G}_{i,j}}\nu_{k}~{}o_{k}~{}T_{k}~{}G_{k}(r_{i},\theta_{j})$ (11)
-where we introduce $\nu_{k}\in[0,1]$ as the acoustic reflectance of the Gaussian and $T_{k}$ as the transmittance from the sensor origin to the Gaussian along the range.
+$I_{u}(r_{i},\theta_{j})=\frac{1}{R_{k}}\sum_{G_{k}^{\prime}\in\mathcal{G}_{i,j}}\nu_{k}~{}o_{k}~{}T_{k}~{}G_{k}^{\prime}(r_{i},\theta_{j})$ (8)
+where $R_{k}$ is the range from the sensor, $T_{k}$ is the transmittance from the sensor origin to the Gaussian along the range, and $G^{\prime}_{k}$ is the Gaussian evaluated in range/azimuth space.
 The subscript indicates that $I_{u}$ is the unsaturated image, meaning that it does not account for the azimuth streaking that causes saturation across the image.
 We observe that the acoustic reflectance of a set of points depends on material properties and viewing angle, so $\nu_{k}$ uses spherical harmonics to encode view-dependent effects.
-Figure 3: Azimuth streaking is characterized by black streaks (referred to by blue arrows) spanning the azimuth axis of a sonar image.
-(a) Sonar image of a metal frame structure that experiences azimuth streaking.
-(b) Sonar image of a concrete piling that has multiple azimuth streaks.
-Image brightness was enhanced for illustrative purposes.
-Images are polar (range/azimuth).
-3.4 Modeling Azimuth Streaking
+In practice, our sonar images are gain adjusted, so $\frac{1}{R_{k}}$ is dropped.
+III-C Azimuth Streak Modeling (ASM)
 Azimuth streaking, as shown in Fig.
-3, is a phenomena that is frequently observed in sonar images.
+4, is a phenomena that is frequently observed in sonar images.
 It is a form of saturation that occurs when the sonar receiver receives strong returns from incident angles close to parallel at a specific range bin $r_{i}$ [32].
-The azimuth streaking model found in [29] proposes to modify all the returns in a range interval if a certain percentage of returns exceeds a given threshold $\zeta$.
-The image with added azimuth streaks is given by:
-$\hat{I}(r_{i},\cdot)=1-(1-I(r_{i},\cdot))^{2}$ (12)
-where $\hat{I}(r_{i},\cdot)$ is the streaked image, and $I(r_{i},\cdot)$ is the vector of intensities at range interval $r_{i}$.
 In our model, we assign each Gaussian a probability that it will cause an azimuth streak.
-During the splatting process, we accumulate this probability as a function of the Gaussian’s opacity, mean, and covariance.
 The benefit of defining this probability per-Gaussian is that it can be optimized to produce azimuth streaks that are multi-view consistent.
-Let $p_{k}\in[0,1]$ be the probability that Gaussian $k$ will contribute to an azimuth streaking phenomena.
-We can then splat these per-Gaussian probabilities into a range-azimuth bin to obtain an azimuth streak probability image $P_{a}$:
-$P_{a}(r_{i},\theta_{j})=\sum_{G_{k}\in\mathcal{G}_{i,j}}p_{k}~{}o_{k}~{}T_{k}~{}G_{k}(r_{i},\theta_{j})$ (13)
+Let $p_{k}$ be the probability that Gaussian $k$ will contribute to an azimuth streaking phenomena.
+We can then splat these per-Gaussian probabilities into a range/azimuth bin to obtain an azimuth streak probability image $P_{a}$:
+$P_{a}(r_{i},\theta_{j})=\sum_{G^{\prime}_{k}\in\mathcal{G}_{i,j}}p_{k}~{}o_{k}~{}T_{k}~{}G^{\prime}_{k}(r_{i},\theta_{j})$ (9)
 $M_{a}(r_{i})=\sum_{\theta=0}^{N_{a}}P_{a}(r_{i},\theta)$ (10)
@@
 The azimuth streak probabilities across the range bins are then used to compute the final image using an adaptive gain mechanism.
-3.4.1 Adaptive Gain Module
+III-C1 Adaptive Gain Module
 We introduce a novel adaptive gain term, $A$, to transform the azimuth streak probability image into receiver gains.
@@
 First, when no Gaussian in a range bin $r_{i}$ has a high probability of azimuth streaking, the gain should be unity.
-Secondly, suppose a single Gaussian $G_{k}$ exhibits a high probability of azimuth streaking across the range bin $r_{i}$.
+Second, suppose a single Gaussian $G_{k}$ exhibits a high probability of azimuth streaking across the range bin $r_{i}$.
 In that case, we wish to adaptively suppress the other Gaussians in $r_{i}$ and assign a high gain to $G_{k}$.
 Finally, if multiple Gaussians in $r_{i}$ have high probabilities of contributing to an azimuth streak, higher gain should be assigned to those Gaussians.
-To address these observations, the adaptive gain term is applied to each range-azimuth bin $(r_{i},\theta_{j})$:
-$A(r_{i},\theta_{j})=P_{a}(r_{i},\theta_{j})M_{a}(r_{i})\frac{e^{\gamma\cdot P_{a}(r_{i},\theta_{j})}+1}{e^{\gamma}+1}+(1-M_{a}(r_{i}))$
+Following these observations, the adaptive gain term is generated for each pixel: $(r_{i},\theta_{j})$:
+$A(r_{i},\theta_{j})=P_{a}(r_{i},\theta_{j})M_{a}(r_{i})\frac{e^{\gamma\cdot P_{a}(r_{i},\theta_{j})}-1}{e^{\gamma}-1}+(1-M_{a}(r_{i}))$
 where $\gamma$ is a scaling factor that dictates the steepness of the adaptive gain.
-This behavior aligns with the insight presented in [29]: if a certain percentage of returns exceeds a threshold, an operation is applied to all the values in a range interval.
+This behavior aligns with the insight presented in [32]: if a certain percentage of returns exceeds a threshold, an operation is applied to all of the values in a range interval.
 However, our proposed model differs because we do not restrict the gain to a fixed function (quadratic) but rather offer a family of curves for our splatting model to explore during optimization.
-Further discussion on the adaptive gain term can be found in the supplementary material.
-The final image with the azimuth streaks is then computed by applying the gain $A(r_{i},\theta_{j})$ to its corresponding bin on $I_{U}(r_{i},\theta_{j})$:
+The final sonar image $\hat{I}(r_{i},\theta_{j})$ with azimuth streaks is then computed by:
 $\hat{I}(r_{i},\theta_{j})=A(r_{i},\theta_{j})\cdot I_{u}(r_{i},\theta_{j})$ (12)
-resulting in the final rendered sonar image $\hat{I}(r_{i},\theta_{j})$.
-Figure 4: We present selected images from our diverse set of 8 sequences used to evaluate of our method.
-We show RGB images of the environment for visualization, and captured range/azimuth sonar images.
+Figure 3: Illustration of the proposed Elevation Sampling Densification Strategy (ESDS).
+We randomly sample $N_{p}$ pixels (shown as green sectors) based on the magnitude of $\mathcal{L}$ then randomly place $N_{g}$ Gaussians (shown as green ellipses) on the elevation arc of that pixel.
+Figure 4: We present our robotic platform and selected images from our diverse set of 8 sequences used in our experiments.
+We show RGB images in top right corners for visualization.
 Our evaluation datasets focus on smaller objects and larger-scale structures such as the test-tank basin.
-Sonar images are polar (range/azimuth).
-3.5 Model Optimization
-We optimize for desired parameters using SGD and by taking losses ($\mathcal{L}_{l1},\mathcal{L}_{ssim}$) between $\hat{I}$ and the ground truth image $I_{\text{GT}}$[13].
+Azimuth streaks clearly occur in the infra_360 and piling_1 sequences.
+Images are range/azimuth.
+III-D Model Optimization
+We optimize for desired parameters using gradient descent and by taking losses ($\mathcal{L}_{l1},\mathcal{L}_{ssim}$) between $\hat{I}$ and the ground truth image $I_{\text{GT}}$ [14].
+Next, we note that large Gaussians cause increased training times and more distortion to the model presented in Sec.
+III-B.
+To prevent Gaussians from becoming too large, we use $\mathcal{L}_{size}=\text{ReLU}(s-s_{init})$, which encourages the scale of the Gaussians $s$ to remain smaller than their initialized scale $s_{init}$ [16].
+Additionally, we find it useful for reconstruction tasks to supervise the opacities in regions with sonar returns via an opacity loss $\mathcal{L}_{o}$, similar to [35].
+We compute $\mathcal{L}_{o}$ between the predicted opacity map $\hat{I}_{o}$ and $I_{\text{GT}}$ thresholded at $\tau_{o}$ to isolate strong returns.
+The threshold parameter is found experimentally for each dataset.
+$\mathcal{L}_{o}$ then becomes an $l1$ loss between $\hat{I}_{o}$ and $I^{o}_{\text{GT}}$.
+The final loss becomes:
+$\mathcal{L}=\lambda_{l1}\mathcal{L}_{l1}+(1-\lambda_{l1})\mathcal{L}_{ssim}+\lambda_{o}\mathcal{L}_{o}+\lambda_{size}\mathcal{L}_{size}$ (13)
+where $\lambda_{l1},\lambda_{o}~{}\text{and}~{}\lambda_{size}$ are the weights for the $l1$, opacity, and size losses respectively.
+III-E Elevation Sampling Densification Strategy (ESDS)
+Splitting and cloning Gaussians in image space is unintuitive for imaging sonar given the elevation angle ambiguity.
+It is possible that a Gaussian can be split then immediately occluded by high opacity Gaussians in front of it.
+To address this, we introduce an alternative strategy for densification of Gaussians called the Elevation Sampling Densification Strategy (ESDS) shown in Fig.
+3.
+Given predicted image $\hat{I}$ and ground truth $I_{\text{GT}}$, we sample $N_{p}$ pixels, where each pixel’s probability is proportional to the magnitude of the loss at that pixel.
+We initialize $N_{g}$ Gaussians on each elevation arc specified by the sampled pixel locations.
+This primarily serves to provide SonarSplat enough opportunities to resolve the elevation angle ambiguity and any possible occlusions by placing Gaussians at different elevation angles.
+As the optimization continues, low-opacity Gaussians are pruned similar to [14].
+III-F Implementation Details
+SonarSplat is built on the gsplat repo [52].
+We train our method on an NVIDIA RTX 3090 GPU with 24 GB of vRAM.
+We calculate the frames per second (FPS) of all methods on the same NVIDIA A6000 GPU with 50 GB of vRAM.
 To properly optimize for the azimuth streaking probabilities $p_{k}$ of each Gaussian, we first identify where azimuth streaks occur by calculating the average intensity of the range interval $r_{i}$.
-Then, for the first $N_{s}$ iterations, we train only on pixels in range bins that exceed a certain average intensity threshold $\tau$.
+Then, for the first $N_{s}$ iterations, we train only on pixels in range bins that exceed a certain average intensity threshold $\tau_{s}$.
+This avoids incorrectly supervising the splat in azimuth streaked regions.
 Importantly, $p_{k}$ is not optimized during this interval.
-Then, after $N_{s}$ iterations, we train only on pixels less than $\tau$.
-We do not optimize $\mu_{k},\Sigma_{k},o_{k}$, or $r_{k}$ during this interval.
+After $N_{s}$ iterations, we optimize only $p_{k}$ in order to capture the azimuth streaking effects.
+We do not optimize $\mu_{k},~{}\Sigma_{k},~{}o_{k}$, or $\nu_{k}$ during this interval.
 This way, we allow the Gaussians to initialize and fit the images before isolating and optimizing the azimuth streaking parameters.
-We also encourage each Gaussian to have either high or low opacity via an opacity loss $\mathcal{L}_{o}$.
-This prevents a set of medium-opacity Gaussians from producing the same pixel intensity.
-Rather, the model is encouraged to use the reflectance parameter $r_{k}$ to represent low-intensity returns.
-We use the negative log-likelihood of two Laplacian distributions as in [15].
-$\mathcal{L}_{o}(x)=|x||1-x|$ (17)
-The final loss becomes:
-$\mathcal{L}=\lambda_{l1}\mathcal{L}_{l1}+\lambda_{ssim}\mathcal{L}_{ssim}+\lambda_{o}\mathcal{L}_{o}$ (18)
-where $\lambda_{l1},\lambda_{ssim},\text{and}~{}\lambda_{o}$ are the weights for the L1, SSIM, and opacity losses respectively.
-4 Results & Experiments
-4.1 Datasets
-We note that real, public datasets of posed sonar images suitable for novel view synthesis experiments do not exist.
-Rather, we collect sequences using our robotic platform based on a BlueRobotics BlueROV2 vehicle shown in Fig.
-1.
-Our robotic platform is equipped with a BluePrint Subsea Oculus M750-d imaging sonar capable of capturing data across various ranges with a 130 horizontal field of view and a 20 vertical field of view.
+IV Results & Experiments
+Figure 5: We present qualitative novel view synthesis results from selected datasets and selected baselines.
+Note that SonarSplat produces more realistic sonar image synthesis compared to baselines and is better able to capture finer details in the environment.
+Note that ZSplat* indicates ZSplat trained only with the sonar loss.
+Images are Cartesian.
+IV-A Robotic Platform and Datasets
+We note that real, public datasets of posed sonar images suitable for novel view synthesis and reconstruction experiments do not exist.
+Rather, we collect sequences using a BlueRobotics BlueROV2 platform, shown in Fig.
+4.
+Our robotic platform is equipped with a BluePrint Subsea Oculus M750-d imaging sonar operating at 1.2MHz, capturing sequences between 3-20m in range, with a 130 horizontal field of view and a 20 vertical field of view.
+We tilt the sonar both at 30 and 0 downwards for various sequences.
 We use this platform to collect 8 trajectories inspecting a range of objects and in both a test tank and river environment.
-Fig.
-4 shows selected sequences for visualization.
-For detailed information about sonar parameters, sensor pose estimation, our platform, and dataset details, please refer to the supplemental material.
-4.2 Baselines
-For baselines, we compare to Neusis [30], Neusis-NGP [44], DSC [7], and ZSplat*.
-Note that ZSplat [32] was originally designed for camera-sonar fusion.
-Since we are focused on sonar-only rendering, we compare to a modified version, ZSplat*, trained with the sonar loss.
-Figure 5: We present qualitative novel view synthesis results from selected datasets and baselines.
-Note that SonarSplat produces superior sonar image synthesis compared to baselines and is better able to capture finer details in the environment.
-SonarSplat is also able to capture azimuth streaking phenomena, which no other baseline is able to do.
-We increase the brightness for all outputs by 25% for easier viewing.
-Note that Z-Splat* indicates Z-Splat trained only with the sonar loss.
-Images are cartesian.
-De-Streaking Performance (ICV) ($\uparrow$)
-Sequence Captured Image SonarSplat NVS De-Streaked Image
-basin_empty 2.74 9.78 19.44
-infra_360 2.89 34.30 44.18
-piling_1 2.78 11.47 19.51
-basin_infra 2.77 4.41 4.84
-rock_360 2.86 25.00 44.47
-piling_2 2.79 7.61 10.32
-pole_river 3.11 2.69 2.78
-rock_river 3.15 5.15 6.78
-Average 2.89 12.55 19.04
-Table 1: Quantitative results for azimuth streak removal. We report the inverse coefficient of variation (ICV) [37]. Higher is better ($\uparrow$), and best is in bold.
-4.3 Novel View Synthesis
-We train our method on an NVIDIA RTX 3090 GPU with 24 GB of vRAM.
-For detailed implementation details, please refer to the supplemental material.
-We report validation PSNR, SSIM, and LPIPS following [13, 21].
+We use 7 sequences for evaluating image synthesis, and 1 sequence with higher coverage of the inspected object for evaluating 3D reconstruction.
+Sensor poses are estimated by filtering a Doppler Velocity Logger and IMU, similar to [1, 33].
+We construct ground truth for 3D reconstruction using RGB images and MASt3R, scaling the reconstruction to metric scale [17].
+We find that MASt3R produces less artifacts than traditional photogrammetry pipelines like AgiSoft Metashape [21].
+IV-B Baselines
+For baselines, we compare to Neusis [33], Neusis-NGP [50], DSC [6], and ZSplat* [35].
+Since we are focused on sonar-only rendering, we compare to a modified version, ZSplat*, trained with only the sonar loss since there is no other work that has explored Gaussian splatting for imaging sonar [35].
+All baselines were trained using official released code and hyperparameters were tuned to the best of our ability.
+TABLE I: Novel view synthesis performance on real world held-out validation data.
+Note that ZSplat* is ZSplat with only the sonar loss activated for training.
+First, second, and third best are shown in red, orange, and yellow, respectively.
+Sonar Image Synthesis
+Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS $\uparrow$
+Neusis [33] 14.68 0.04 0.64 0.03
+Neusis-NGP [50] 19.56 0.33 0.64 0.01
+DSC [6] 14.70 0.07 0.66 0.03
+ZSplat* [35] 10.06 0.02 0.69 155.32
+SonarSplat (ours) 22.79 0.41 0.48 13.45
+IV-C Novel View Synthesis
+We report validation PSNR, SSIM, and LPIPS following [14], using the lpipsPyTorch library.
 For our validation set, we take every 8th image from the dataset as introduced in [26].
 We report SonarSplat’s performance compared to baselines in Tab.
-2.
-We find that SonarSplat consistently synthesizes more realistic sonar images, quantified by its higher PSNR, SSIM, and lower LPIPS values.
-We also present qualitative results from selected sequences in Fig.
+I.
+We find that SonarSplat consistently synthesizes more realistic sonar images, quantified by its higher PSNR and SSIM, and lower LPIPS values.
+Qualitative results from selected sequences are presented in Fig.
 5.
-Note that these qualitative results are in cartesian space for interpretability.
-For extensive polar image results, please refer to the supplemental material.
-Finally, we explore the feasibility of SonarSplat for data synthesis applications by reporting the rendering speed in frames per second (FPS) of each method in Tab.
-3.
-We calculate the FPS of each method on an NVIDIA A6000 GPU with 50 GB of vRAM.
-SonarSplat and ZSplat both leverage Gaussian splatting, which offers significantly faster rendering speeds compared to neural rendering methods.
-ZSplat is optimized for high rendering speed, which will be a focus of future work for SonarSplat.
-Sonar Image Synthesis
-Method PSNR ($\uparrow$) SSIM ($\uparrow$) LPIPS ($\downarrow$)
-Neusis [30] 14.43 0.03 0.66
-Neusis-NGP [44] 19.43 0.33 0.55
-DSC [7] 14.87 0.10 0.63
-Z-Splat* [32] 8.57 0.03 0.68
-SonarSplat (ours) 21.89 0.41 0.62
-Table 2: Novel view synthesis performance on real world data. Note that ZSplat* is ZSplat with only the sonar loss activated for training. Best is shown in bold.
-Figure 6: Qualitative results for azimuth streak removal capabilities of SonarSplat.
-Note that other baselines are unable to remove azimuth streaks from synthesized images since they do not learn the streaking parameters.
-Locations of azimuth streaks in the captured images are shown in blue arrows for illustrative purposes.
-SonarSplat is able to effectively restore sonar imagery that exhibits azimuth streaking.
-Images are cartesian.
-Rendering Speed ($\uparrow$)
-Method FPS
-Neusis [30] 0.03
-Neusis-NGP [44] 0.01
-DSC [7] 0.03
-Z-Splat* [32] 155.32
-SonarSplat (ours) 9.46
-Table 3: Rendering speed in frames per second (FPS) averaged across all validation images across all sequences.
-All results are reported on an NVIDIA RTX A6000 GPU.
-Note that ZSplat* is ZSplat with only sonar rendering.
-Best is shown in bold.
-Figure 7: Qualitative results from 3D reconstruction of SonarSplat compared to baselines.
-4.4 Azimuth Streak Removal
-Since SonarSplat learns probabilities of azimuth streaking for each Gaussian, we can undo the adaptive gain to render images without azimuth streaks present and recover suppressed returns.
+We also report the rendering speed in FPS of each method in Tab.
+I.
+SonarSplat and ZSplat* both leverage Gaussian splatting, which offers significantly faster rendering speeds compared to neural rendering methods.
+ZSplat* is optimized for high rendering speed, which will be a focus of future work.
+SonarSplat Rendering Ablations
+VDR
+ASM
+ESDS
+PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+21.32 0.35 0.53
+21.16 0.34 0.55
+22.51 0.40 0.52
+22.79 0.41 0.48
+TABLE II: Ablation study results for the proposed acoustic reflectance parameter (VDR), azimuth streak modeling (ASM), and elevation sampling densification strategy (ESDS). Best in bold.
+IV-D Ablation Studies
+We perform ablation studies on the view-dependent reflectivity parameter $\nu_{k}$ (VDR), the azimuth streak modeling (ASM), and the Elevation Sampling Densification Strategy (ESDS).
+Ablation studies are averaged across 7 evaluation datasets and trained with exactly the same hyperparameters as results from Sec.
+IV-C.
+Methods that don’t use ESDS use the mcmc strategy in gsplat [52].
+Results are reported in Tab.
+II and shown in Fig.
+6.
+Clearly, the addition of azimuth streak modeling, view-dependent acoustic reflectivity, and the elevation sampling densification strategy improves the image synthesis capabilities of SonarSplat.
+Figure 6: Qualitative comparison of SonarSplat without and with azimuth streak modeling (ASM) for piling_1.
+Note that if ASM is not modeled, rendered views will produce erroneous streaks (indicated with red arrows) instead of capturing the true streaks (indicated with blue arrows).
+Images are polar.
+Figure 7: Qualitative results for azimuth streak removal capabilities of SonarSplat. Locations of azimuth streaks in the captured images are indicated by blue arrows.
+TABLE III: Quantitative results for azimuth streak removal. We report the inverse coefficient of variation (ICV). Higher is better ($\uparrow$) and best is in bold.
+De-Streaking Performance (ICV) $\uparrow$
+Captured Image De-Streaked Prediction
+Average 2.91 18.58
+IV-E Azimuth Streak Removal/De-Streaking
+Using SonarSplat, we can undo the adaptive gain to render images without azimuth streaks present and recover suppressed returns.
 This vision task is most similar to image de-striping from the remote sensing literature.
 We present quantitative results in Tab.
-1 using the inverse coefficient of variation (ICV) metric, which is a common no-reference metric used in de-striping works [36, 33, 26, 37].
+III using the inverse coefficient of variation (ICV) metric, which is a common no-reference metric used in de-striping works [42, 36, 29].
 We present qualitative results of de-streaked images in Fig.
-6.
-4.5 3D Reconstruction
-We show that SonarSplat learns the geometric properties of the scene by converting the optimized splat into a pointcloud for visualization.
-We compare with methods focusing on 3D reconstruction: Neusis and DSC [30, 7].
+7.
+The Azimuth Streak Modeling allows SonarSplat to qualitatively and quantitatively remove streaks effectively and improves the finer features in sonar images.
+Figure 8: Qualitative results from 3D reconstruction of SonarSplat compared to baselines. We inspect a concrete piling in a test tank. Ground truth is created using collected RGB images.
+IV-F Real-World 3D Reconstruction
+3D reconstruction is crucial for inspection and visualization of underwater structures. In a test tank, we demonstrate inspection of a concrete piling, a structure commonly found in bridges and dams.
+We show that SonarSplat learns the geometric properties of the scene by converting the optimized splat into a mesh for visualization.
+We compare to state-of-the-art sonar reconstruction methods: Neusis, DSC, and ZSplat* [6, 33, 35].
 Qualitative results are shown in Fig.
 8 and quantitative results are shown in Tab.
-4.
-We use popular metrics from scene reconstruction literature and choose $l1$ Chamfer Distance, F-Score, and voxelized IoU to evaluate performance [1].
-For Neusis and DSC, we produce meshes using marching cubes.
-For SonarSplat, we choose to sample points from the 3D Gaussians to produce a pointcloud.
-We sample 100,000 points from both the mesh and Gaussian splat and compute evaluation metrics against a ground truth photogrammetry reconstruction of the rock platform.
-We note that neither Neusis nor DSC are able to reconstruct the object whereas SonarSplat is able to produce a coarse representation.
+IV.
+Following [33, 6], we report Root Mean Squared (RMS) $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD).
+We produce meshes from all methods then randomly sample 30,000 points from the surface for 30 iterations to calculate RMS.
+Note that neither DSC, Neusis, nor ZSplat* is able to reconstruct the object whereas SonarSplat is able to produce an accurate reconstruction of the piling.
+TABLE IV: Quantitative evaluation of 3D reconstruction performance on the piling using RMS $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD). Best in bold.
 Reconstruction Metrics
-Method CD-$l_{1}$ (m) $\downarrow$ F-Score $\uparrow$ IoU $\uparrow$
-Neusis [30] 0.202 0.4834 0.0129
-DSC [7] 0.187 0.4613 0.0164
-SonarSplat (ours) 0.084 0.1170 0.0191
-Table 4: We quantitatively evaluate the 3D reconstruction of selected methods on the rock_360 sequence. Best is in bold.
-5 Conclusion & Future Work
-We propose SonarSplat, the first sonar-only Gaussian splatting framework for novel view synthesis for imaging sonar in underwater applications.
-First, we adapt the sonar rendering equation for efficient range-azimuth splatting by evaluating 3D Gaussians.
+Method CD-$l_{1}$downarrow HD $\downarrow$
+Neusis [33] 0.32 0.78
+DSC [6] 0.29 0.82
+ZSplat* [35] 2.93 4.88
+SonarSplat (ours) 0.14 0.57
+V Conclusion & Future Work
+We propose SonarSplat, the first sonar-only Gaussian splatting framework for imaging sonar in underwater applications.
+First, we adapt the sonar rendering equation for efficient range/azimuth splatting by evaluating 3D Gaussians.
 We model azimuth streaking within the Gaussian splatting framework, allowing us to learn per-Gaussian azimuth streaking probabilities and produce high-quality de-streaked images.
-We perform experiments on real-world data in test tank and river environments.
-In the novel view synthesis task, SonarSplat outperforms baselines by +2.5 dB PSNR and demonstrates superior qualitative results.
-We also demonstrate that SonarSplat learns scene geometry, demonstrated by both quantitative and qualitative results.
+We also introduce a novel densification strategy (ESDS) to enable SonarSplat to resolve elevation angle ambiguities and associated occlusions.
+From our extensive experiments on a diverse set of real-world data in test tank and river environments, we find that SonarSplat outperforms state-of-the-art methods in novel view synthesis (+3.2 dB PSNR) and 3D reconstruction (52% lower Chamfer Distance), and additionally allows for de-streaking of sonar images.
 Future work will focus on leveraging SonarSplat for sonar data synthesis by randomizing scene parameters including acoustic reflectance and azimuth streaking probability.
-This can yield more diverse datasets for training sonar-based scene understanding algorithms.
+This can yield more diverse datasets for training sonar-based scene understanding algorithms, including multi-view target recognition algorithms [40, 38].
 We will also work to improve SonarSplat’s ability to perform dense 3D reconstruction by better modeling additional acoustic phenomena like speckle noise, Lambertian scattering, and multi-path reflections.
-6 Acknowledgment
-This work is supported by the National Science Foundation under Award No. 2337774.
+Our choice of Gaussian splatting as a 3D scene representation also opens possibilities for real-time sensor fusion and SLAM [13, 23].

### v3

@@
 We evaluate SonarSplat using real-world datasets of sonar images collected from an underwater robotic platform in a controlled test tank and in a real-world river environment.
-Compared to the state-of-the-art, SonarSplat offers improved image synthesis capabilities (+3.2 dB PSNR) and more accurate 3D reconstruction (52% lower Chamfer Distance).
+Compared to the state-of-the-art, SonarSplat offers improved image synthesis capabilities (+3.2 dB PSNR) and more accurate 3D reconstruction (77% lower Chamfer Distance).
 We also demonstrate that SonarSplat can be leveraged for azimuth streak removal.
@@
 Acoustic sensors, such as imaging sonar, are commonly used for infrastructure inspection, large-area mapping, and target detection in underwater environments [1, 2, 3].
-Unlike optical sensors, which are severely range-limited due to water column effects on light propagation, acoustic sensors can capture data at long ranges to provide critical information about subsea environments.
-Although sonar exhibits many desirable qualities such as longer range, invariance to lighting conditions, and the ability to discern material properties, there exist acoustic phenomena, including elevation ambiguity, azimuth streaking, and multi-path reflections, which make sonar interpretation difficult for operators and computer vision algorithms alike.
+While optical sensors are severely range-limited due to water column effects on light propagation, acoustic sensors can capture data at long ranges to provide critical information about subsea environments.
+However, acoustic phenomena like elevation ambiguity, azimuth streaking, and multi-path reflections make sonar interpretation difficult for both operators and computer vision algorithms.
 Figure 1: We present SonarSplat, a novel 3D Gaussian splatting framework that enables novel view synthesis of realistic sonar images, image de-streaking, and 3D reconstruction.
@@
 Most relevant to our proposed work, ZSplat proposes a Gaussian splatting framework for RGB-sonar fusion [19].
-However, ZSplat leverages the fusion of sonar data to improve the rendering of RGB images and does not focus on enabling high-fidelity data synthesis for sonar imagery or provide evaluation for the quality of rendered sonar images or 3D reconstruction.
+However, ZSplat leverages the fusion of sonar data to improve the rendering of RGB images and does not focus on enabling high-fidelity data synthesis for sonar imagery or providing evaluation for the quality of rendered sonar images or 3D reconstruction.
 Thus, there is a clear gap in the literature for a framework capable of efficient and effective sonar image synthesis and 3D reconstruction using Gaussian splatting.
 In this paper, we propose SonarSplat (Fig.
-1), a novel Gaussian splatting framework for imaging sonar that enables efficient and high quality novel view synthesis, sonar image de-streaking, and 3D reconstruction for underwater robotic applications.
+1), a novel Gaussian splatting framework for imaging sonar that enables efficient and high-quality novel view synthesis, sonar image de-streaking, and 3D reconstruction for underwater robotic applications.
 To the best of our knowledge, SonarSplat is the first sonar-only method that leverages Gaussian splatting to perform these tasks.
-Our main contributions are summarized as follows:
-  • We develop a novel 3D Gaussian splatting framework for rasterization of range/azimuth sonar images from known and novel viewpoints.
+Our main contributions are as follows:
+  • We develop a novel 3D Gaussian splatting framework for rasterization of range/azimuth sonar images from known and arbitrary viewpoints.
   • We develop a novel method for learning the probability of azimuth streaking in a differentiable manner, which allows for removal of unwanted azimuth streaking artifacts from rendered images.
@@
   • We perform quantitative and qualitative evaluation of our method on real imaging sonar data collected from a robotic platform with a variety of sonar ranges and environments.
-Code and data will be made publicly available on a project website after review.
+For further information, please visit our project website at https://umfieldrobotics.github.io/sonarsplat3D/.
 II Related Work
@@
 Additionally, we optimize and splat per-Gaussian azimuth streaking probabilities $p_{k}$.
-All of the probabilities in a range bin $r_{i}$ are considered in our novel Adaptive Gain Module, which adjusts the receiver gain applied to $I_{u}$.
-Finally, we produce $\hat{I}$ by multiplying the gain $A$ by $I_{u}$.
+All of the probabilities in a range bin $r_{i}$ are considered in our novel Adaptive Gain Module, which adjusts the gain AA applied to $I_{u}$.
+Finally, we produce $\hat{I}$ by multiplying the gain AA by $I_{u}$.
 All parameters are optimized using gradient descent by taking losses with respect to the sonar image pixel values.
@@
 II-A Radiance Field Methods
-NeRFs represent scenes implicitly using neural networks and have been explored for a variety of robotic applications, including novel view synthesis, data generation, Simultaneous Localization and Mapping (SLAM), manipulation, image dehazing, and trajectory optimization [26, 2, 11, 37, 43, 5, 7].
+NeRFs represent scenes implicitly using neural networks and have been explored for a variety of robotic applications, including novel view synthesis [4], data generation [20], Simultaneous Localization and Mapping (SLAM) [21, 22], manipulation [23], and image dehazing [7].
 Despite their success as a scene representation, a key challenge for the practical application of NeRFs is the long training time and the slow rendering speeds, primarily due to the computationally expensive volumetric rendering process [4].
-A recent development in radiance field methods is Gaussian splatting, which uses more efficient rasterization instead of volumetric rendering to optimize and produce a scene representation, enabling usage in a variety of real-time applications [14, 23, 22, 25].
+A recent development in radiance field methods is Gaussian splatting, which uses more efficient rasterization instead of volumetric rendering to produce a scene representation.
+This has led to the use of radiance field methods in a variety of real-time applications [15, 24, 25].
 Despite their impressive performance and rapid adoption, radiance field methods were developed with optical sensors in mind, limiting their direct application to non-visual sensors.
-Recent developments have integrated the physics-based image formation models of various non-visual sensors, including radar, LIDAR, sonar, thermal cameras, and event cameras into NeRF and Gaussian splatting frameworks for various vision tasks [3, 4, 9, 33, 20].
+Recent developments have integrated non-visual sensors, including radar, LIDAR, sonar, thermal cameras, and event cameras into NeRF and Gaussian splatting frameworks for various vision tasks [26, 27, 21, 28, 10, 29].
 We are inspired by the faster rendering speeds and high fidelity of Gaussian splatting methods and propose a novel Gaussian splatting framework for imaging sonar.
@@
 However, the sensor is susceptible to multi-path effects, nonlinear noise patterns, and phenomena like azimuth streaking [33].
-Azimuth streaking, in particular, can make sonar image understanding difficult by suppressing returns from certain azimuth angles, causing black bars in the sonar image [31].
+Azimuth streaking, in particular, can make sonar image understanding difficult by suppressing returns from specific ranges, causing black bars in the sonar image [33].
 In addition, the sensor inherently loses information from 3D into 2D when capturing the acoustic returns, causing an effect known as elevation angle ambiguity [34].
@@
 Differentiable Space Carving (DSC) proposes a faster method that models echo probabilities to carve out a mesh of the scene [13].
-Through our experiments, we show that none of these method are able to perform realistic looking novel view synthesis for imaging sonar, limiting their application to data generation and augmentation.
-Furthermore, querying these methods is time consuming and limits their application to real-time methods.
+Through our experiments, we demonstrate that these methods render less realistic sonar imagery compared to SonarSplat according to quantitative metrics including PSNR, SSIM, and LPIPS.
+This limits their use in data generation and augmentation.
+Additionally, our experiments indicate that querying these methods is time consuming, which limits their use in real-time applications.
 To the best of our knowledge, ZSplat is the only Gaussian splatting paper that incorporates imaging sonar data, with demonstrations on underwater data.
-However, ZSplat oversimplifies the splatting process by not accounting for any sonar-specific phenomena.
+However, ZSplat operates with a simplified sonar model that does not model several important sonar-specific phenomena.
 More importantly, ZSplat only splats the opacities of Gaussians and does not consider material properties like acoustic reflectance.
-This leads to poor sonar renderings, which limits its application to sonar novel view synthesis.
+This leads to poor sonar renderings, which limits its application in sonar novel view synthesis.
 We demonstrate that SonarSplat, which more accurately models sonar image formation, produces more realistic novel view synthesis, restores sonar images to remove azimuth streaking artifacts, and produces 3D reconstructions under a unified framework.
@@
 Figure 2 shows an overview of SonarSplat, which uses sonar images and their corresponding sensor poses to optimize a scene representation.
-The scene is represented as a set of 3D Gaussian primitives, each parametrized with a mean, covariance, opacity, acoustic reflectivity and azimuth streaking probability.
+The scene is represented as a set of 3D Gaussian primitives, each parametrized with a mean, covariance, opacity, acoustic reflectivity, and azimuth streaking probability.
 First, the Gaussians that fall within the view frustum of the sonar are rasterized into a range/azimuth image.
@@
 Finally, we produce a sonar image with azimuth streaks that can be optimized by taking a loss with respect to the ground truth sonar image.
-We provide further detail of each of these components in the remainder of this section.
+We provide further details of each of these components in the remainder of this section.
 III-A Imaging Sonar Formation Model
-To adapt the optical Gaussian splatting framework to accommodate imaging sonar, we must incorporate the acoustic image formation model.
-Imaging sonar is a time-of-flight sensor that can provide the echo intensities at a given range and azimuth bin.
-Importantly, imaging sonar is not able to distinguish the elevation angle of a return, leading to an elevation angle ambiguity.
-This ambiguity makes direct use of imaging sonar difficult for 3D reconstruction and triangulation tasks [47, 10].
+To adapt the optical Gaussian splatting framework for imaging sonar, we must incorporate the acoustic image formation model into the rasterization process.
+Imaging sonar is a time-of-flight sensor that provides the echo intensities at a given range and azimuth bin.
 A sonar image is defined as $I\in\mathbb{R}^{N_{a}\times N_{r}}$ where $N_{a}$ denotes the number of azimuth bins and $N_{r}$ denotes the number of range bins.
 In spherical coordinates, the sonar has a vertical field of view ($\phi_{min},\phi_{max}$), a horizontal field of view ($\theta_{min},\theta_{max}$), and a maximum range, $R_{max}$.
-Finally, taking $\frac{R_{max}}{N_{r}}$ gives us the range resolution, $\epsilon_{r}$, and taking $\frac{\theta_{max}-\theta{min}}{N_{a}}$ gives us the azimuthal resolution, $\epsilon_{a}$.
+Finally, $\frac{R_{max}}{N_{r}}$ is the range resolution, $\epsilon_{r}$, and $\frac{\theta_{max}-\theta_{min}}{N_{a}}$ is the azimuthal resolution, $\epsilon_{a}$.
 We leverage the sonar rendering equation introduced in [10], which calculates the intensity of return at bin $(r_{i},\theta_{j})$ in the sonar image $I$ by:
-[…]
+$I(r_{i},\theta_{j})=\int_{\phi_{min}}^{\phi_{max}}\int_{\theta_{j}-\epsilon_{a}}^{\theta_{j}+\epsilon_{a}}\int_{r_{i}-\epsilon_{r}}^{r_{i}+\epsilon_{r}}\frac{E_{e}}{r}T(r,\theta,\phi)\sigma(r,\theta,\phi)drd\theta d\phi$
 where $E_{e}$ is the intensity of emitted sound, $r$ is the range of the point, $T(r,\theta,\phi)$ is the transmittance term, and $\sigma(r,\theta,\phi)$ is the density of the point.
-Note that this transmittance term is equivalent to the term used in volumetric rendering, serving as the probability that a ray travels to point $(r,\theta,\phi)$ without hitting another particle.
+$T(r,\theta,\phi)$ is the probability that a ray travels to point $(r,\theta,\phi)$ without hitting another particle.
 III-B Range/Azimuth Splatting Model
@@
 To this end, we represent the scene with a set of 3D Gaussian primitives $\mathcal{G}=\{G_{k}\}_{k=1}^{N}$.
-We augment this parametrization with acoustic-specific variables: acoustic reflectivity $\nu_{k}$ and azimuth streak probability $p_{k}$.
-Formally, these Gaussians are parametrized with a mean $\mu_{k}\in\mathbb{R}^{3}$, scale $s_{k}\in\mathbb{R}^{3}$, orientation $q_{k}\in\mathbb{H}$, opacity $o_{k}\in\mathbb{R}^{+}$, acoustic reflectivity $\nu_{k}\in\mathbb{R}^{+}$, and azimuth streak probability $p_{k}\in\mathbb{R}^{+}$:
-$G_{k}=\{\mu_{k},s_{k},q_{k},o_{k},\nu_{k},p_{k}\}$ (2)
-where the covariance $\Sigma_{k}$ is computed as in [14].
+Formally, these Gaussians are parametrized with a mean $\mu_{k}\in\mathbb{R}^{3}$, scale $s_{k}\in\mathbb{R}^{3}$, orientation $q_{k}\in\mathbb{H}$, opacity $o_{k}\in\mathbb{R}^{+}$, and azimuth streak probability $p_{k}\in\mathbb{R}^{+}$.
+We take inspiration from other Time-of-Flight novel view synthesis work and use spherical harmonics coefficients $sh_{k}$ to model view-dependent acoustic reflectivity $\nu_{k}$ [28].
+Specifically, $\nu_{k}=SH(sh_{k},\vec{v})$, where $\vec{v}$ is the viewing direction.
+$G_{k}=(\mu_{k},s_{k},q_{k},o_{k},sh_{k},p_{k})$ (2)
+where the covariance $\Sigma_{k}$ $R_{k}S_{k}S_{k}^{T}R_{k}^{T}$ for rotation matrix $R_{k}$ given by quaternion $q_{k}$ and scaling matrix $S_{k}$ given by $s_{k}$.
 To render sonar images, we first transform the means and covariances of the Gaussians into spherical coordinates.
-Let $\mu_{k}=\langle x_{k},y_{k},z_{k}\rangle~{}\text{and}~{}\Sigma_{k}$ be the mean and covariance in the sensor’s frame, respectively.
+Let $\mu_{k}=[x_{k},y_{k},z_{k}]^{T}~\text{and}~\Sigma_{k}$ be the mean and covariance in the sensor’s frame, respectively.
 The transformation to spherical coordinates is as follows [37]:
-$\mu_{s}=\begin{bmatrix}r_{k}\\ \theta_{k}\\ \phi_{k}\end{bmatrix}=\begin{bmatrix}||\mu_{k}||\\ \arctan(y_{k},x_{k})\\ \arctan(z_{k},\sqrt{x_{k}^{2}+y_{k}^{2}})\\ \end{bmatrix}$ (3)
+[…]
 The covariance matrix $\Sigma_{k}$ is transformed by using a first-order linearization of the coordinate transformation, which the Jacobian describes:
-$J_{s}=\begin{bmatrix}\frac{x_{k}}{\lVert\mu_{k}\rVert}&\frac{y_{k}}{\lVert\mu_{k}\rVert}&\frac{z_{k}}{\lVert\mu_{k}\rVert}\\[4.30554pt] \frac{-y}{x^{2}+y^{2}}&\frac{x}{x^{2}+y^{2}}&0\\[4.30554pt] \frac{-x_{k}\cdot z_{k}}{||\mu_{k}||^{2}\sqrt{x^{2}+y^{2}}}&\frac{-y_{k}\cdot z_{k}}{||\mu_{k}||^{2}\sqrt{x^{2}+y^{2}}}&\frac{\sqrt{x^{2}+y^{2}}}{\lVert\mu_{k}\rVert^{2}}\end{bmatrix}$ (4)
+[…]
 and the covariance matrix becomes
@@
 $K=\begin{bmatrix}\frac{1}{\epsilon_{r}}&0&0\\ 0&\frac{1}{\epsilon_{a}}&\frac{N_{a}}{2}\\ 0&0&1\\ \end{bmatrix}$ (6)
-and then we perform the transformation
+and then we perform the transformation:
 $\mu_{k}^{\prime}=K\mu_{s},\ \ \Sigma_{k}^{\prime}=J_{K}W\Sigma_{s}WJ_{K}^{T}$ (7)
 where $\mu_{k}^{\prime}$ and $\Sigma_{k}^{\prime}$ are the mean and covariance transformed to range/azimuth space, $J_{K}$ is the Jacobian of $K$, and $W$ is the sensor’s view matrix, similar to the camera splatting formulation [15].
-To find the intensity of acoustic returns at a given range/azimuth bin ($r_{i},\theta_{j}$), we consider $\mathcal{G}_{i,j}\subset\mathcal{G}$, the set of Gaussians transformed to range/azimuth image space that overlap with the range/azimuth bin ($r_{i},\theta_{j}$).
+To find the intensity of acoustic returns at a given range/azimuth bin $r_{i},\theta_{j}]$, we consider $\mathcal{G}_{i,j}\subset\mathcal{G}$, the set of Gaussians transformed to range/azimuth image space that overlap with the range/azimuth bin $r_{i},\theta_{j}]$.
 The rasterization equation to obtain the sonar image for SonarSplat is similar to that of [15], with a few key differences.
+Specifically, we use single-channel acoustic reflectance $\nu$ instead of per-channel color $c$ and we form the image by splatting in the range/azimuth plane using $r,\theta]$ coordinates instead of traditional camera coordinates.
 SonarSplat’s rasterization equation is given by:
-$I_{u}(r_{i},\theta_{j})=\frac{1}{R_{k}}\sum_{G_{k}^{\prime}\in\mathcal{G}_{i,j}}\nu_{k}~{}o_{k}~{}T_{k}~{}G_{k}^{\prime}(r_{i},\theta_{j})$ (8)
-where $R_{k}$ is the range from the sensor, $T_{k}$ is the transmittance from the sensor origin to the Gaussian along the range, and $G^{\prime}_{k}$ is the Gaussian evaluated in range/azimuth space.
+$I_{u}(r_{i},\theta_{j})=\sum_{G_{k}\in\mathcal{G}_{i,j}}\frac{1}{r_{k}}\nu_{k}~o_{k}~\overline{T}_{k}~f_{k}([r_{i},\theta_{j}])$ (8)
+$f_{k}(x)=\text{exp}(-\frac{1}{2}\lVert x-\mu^{\prime}_{k}\rVert^{2}_{\Sigma^{\prime}_{k}})$ (9)
+$\overline{T}_{k}=\frac{1}{|\mathcal{P}_{k}|}\sum_{p\in\mathcal{P}_{k}}T^{p}_{k}$ (10)
+where $r_{k}$ is the range of $G_{k}$ from the sensor, $f_{k}(x)$ is the evaluation of Gaussian $G_{k}$ at range/azimuth point $r_{i},\theta_{j}]$ with the norm as the Mahalanobis distance, $\overline{T}_{k}$ is the average transmittance of the Gaussian from the sensor origin to the Gaussian along the range.
+To compute $\overline{T}_{k}$, 3D Gaussians are splatted onto the azimuth/elevation plane to model occlusions in the range dimension.
+We define $\mathcal{P}_{k}$ as the set of azimuth/elevation pixels that the Gaussian $G_{k}$ overlaps and we calculate $T^{p}_{k}$ as the transmittance of $G_{k}$ at pixel $p$ according to [15].
 The subscript indicates that $I_{u}$ is the unsaturated image, meaning that it does not account for the azimuth streaking that causes saturation across the image.
-We observe that the acoustic reflectance of a set of points depends on material properties and viewing angle, so $\nu_{k}$ uses spherical harmonics to encode view-dependent effects.
-In practice, our sonar images are gain adjusted, so $\frac{1}{R_{k}}$ is dropped.
+Our sonar images are gain adjusted, so $\frac{1}{r_{k}}$ is dropped.
 III-C Azimuth Streak Modeling (ASM)
@@
 4, is a phenomena that is frequently observed in sonar images.
-It is a form of saturation that occurs when the sonar receiver receives strong returns from incident angles close to parallel at a specific range bin $r_{i}$ [32].
+It is a form of saturation that occurs when the sonar receiver receives strong returns from incident angles close to 0 at a specific range bin $r_{i}$ [33].
 In our model, we assign each Gaussian a probability that it will cause an azimuth streak.
@@
 We can then splat these per-Gaussian probabilities into a range/azimuth bin to obtain an azimuth streak probability image $P_{a}$:
-$P_{a}(r_{i},\theta_{j})=\sum_{G^{\prime}_{k}\in\mathcal{G}_{i,j}}p_{k}~{}o_{k}~{}T_{k}~{}G^{\prime}_{k}(r_{i},\theta_{j})$ (9)
+$P_{a}(r_{i},\theta_{j})=\sum_{G_{k}\in\mathcal{G}_{i,j}}p_{k}~o_{k}~\overline{T}_{k}~f_{k}([r_{i},\theta_{j}])$ (11)
 $M_{a}(r_{i})=\sum_{\theta=0}^{N_{a}}P_{a}(r_{i},\theta)$ (12)
@@
 III-C1 Adaptive Gain Module
-We introduce a novel adaptive gain term, $A$, to transform the azimuth streak probability image into receiver gains.
+We introduce a novel adaptive gain term, $A$, that is produced by transforming the azimuth streak probability image.
 We begin designing this gain term with a few observations.
@@
 We randomly sample $N_{p}$ pixels (shown as green sectors) based on the magnitude of $\mathcal{L}$ then randomly place $N_{g}$ Gaussians (shown as green ellipses) on the elevation arc of that pixel.
-Figure 4: We present our robotic platform and selected images from our diverse set of 8 sequences used in our experiments.
+Figure 4: We present our robotic platform and selected images from our diverse set of 9 sequences used in our experiments.
 We show RGB images in top right corners for visualization.
@@
 Azimuth streaks clearly occur in the infra_360 and piling_1 sequences.
-Images are range/azimuth.
 III-D Model Optimization
-We optimize for desired parameters using gradient descent and by taking losses ($\mathcal{L}_{l1},\mathcal{L}_{ssim}$) between $\hat{I}$ and the ground truth image $I_{\text{GT}}$ [14].
+We optimize SonarSplat using gradient descent and by taking losses ($\mathcal{L}_{l1},\mathcal{L}_{ssim}$) between $\hat{I}$ and the ground truth image $I_{\text{GT}}$ [15].
 Next, we note that large Gaussians cause increased training times and more distortion to the model presented in Sec.
@@
 To prevent Gaussians from becoming too large, we use $\mathcal{L}_{size}=\text{ReLU}(s-s_{init})$, which encourages the scale of the Gaussians $s$ to remain smaller than their initialized scale $s_{init}$ [38].
-Additionally, we find it useful for reconstruction tasks to supervise the opacities in regions with sonar returns via an opacity loss $\mathcal{L}_{o}$, similar to [35].
+Additionally, we find it useful for reconstruction tasks to supervise the opacities of Gaussians via an opacity loss $\mathcal{L}_{o}$, similar to [19].
 We compute $\mathcal{L}_{o}$ between the predicted opacity map $\hat{I}_{o}$ and $I_{\text{GT}}$ thresholded at $\tau_{o}$ to isolate strong returns.
@@
 $\mathcal{L}=\lambda_{l1}\mathcal{L}_{l1}+(1-\lambda_{l1})\mathcal{L}_{ssim}+\lambda_{o}\mathcal{L}_{o}+\lambda_{size}\mathcal{L}_{size}$ (15)
-where $\lambda_{l1},\lambda_{o}~{}\text{and}~{}\lambda_{size}$ are the weights for the $l1$, opacity, and size losses respectively.
+where $\lambda_{l1},\lambda_{o},~\text{and}~\lambda_{size}$ are the weights for the $l1$, opacity, and size losses respectively.
 III-E Elevation Sampling Densification Strategy (ESDS)
@@
 This primarily serves to provide SonarSplat enough opportunities to resolve the elevation angle ambiguity and any possible occlusions by placing Gaussians at different elevation angles.
-As the optimization continues, low-opacity Gaussians are pruned similar to [14].
+Low-opacity Gaussians are pruned similar to [15].
 III-F Implementation Details
@@
 We calculate the frames per second (FPS) of all methods on the same NVIDIA A6000 GPU with 50 GB of vRAM.
+For every pixel in a sonar image with intensity above a threshold $\tau_{init}$, we initialize $N_{init}$ Gaussians on the elevation arc specified by the pixel’s range, azimuth, and the sensor’s vertical field of view.
+This process is repeated for all sonar images in a dataset to produce the initial set of 3D Gaussians.
 To properly optimize for the azimuth streaking probabilities $p_{k}$ of each Gaussian, we first identify where azimuth streaks occur by calculating the average intensity of the range interval $r_{i}$.
@@
 We do not optimize $\mu_{k},~\Sigma_{k},~o_{k}$, or $\nu_{k}$ during this interval.
-This way, we allow the Gaussians to initialize and fit the images before isolating and optimizing the azimuth streaking parameters.
+This allows the Gaussians to fit the images before optimizing the azimuth streaking parameters.
 IV Results & Experiments
@@
 Our robotic platform is equipped with a BluePrint Subsea Oculus M750-d imaging sonar operating at 1.2MHz, capturing sequences between 3-20m in range, with a 130 horizontal field of view and a 20 vertical field of view.
+The Oculus M750-d has a range resolution varying from 0.5-4 cm/pixel and angular resolution of 0.6 with a beam separation of 0.25 [40].
 We tilt the sonar both at 30 and 0 downwards for various sequences.
-We use this platform to collect 8 trajectories inspecting a range of objects and in both a test tank and river environment.
-We use 7 sequences for evaluating image synthesis, and 1 sequence with higher coverage of the inspected object for evaluating 3D reconstruction.
+We use this platform to collect 9 trajectories inspecting a range of objects and in both a test tank and river environment.
+We use 7 sequences for evaluating image synthesis, and 2 sequences with higher coverage of the inspected object for evaluating 3D reconstruction.
+Our 3D reconstruction datasets are called Concrete Piling and Monohansett Shipwreck.
 Sensor poses are estimated by filtering a Doppler Velocity Logger and IMU, similar to [41, 10].
-We construct ground truth for 3D reconstruction using RGB images and MASt3R, scaling the reconstruction to metric scale [17].
-We find that MASt3R produces less artifacts than traditional photogrammetry pipelines like AgiSoft Metashape [21].
+For Concrete Piling, we construct ground truth for 3D reconstruction using RGB images and MASt3R, scaling the reconstruction to metric scale [42].
+For Monohansett Shipwreck, we use an underwater LiDAR scan of the shipwreck as ground truth.
 IV-B Baselines
@@
 Since we are focused on sonar-only rendering, we compare to a modified version, ZSplat*, trained with only the sonar loss since there is no other work that has explored Gaussian splatting for imaging sonar [19].
-All baselines were trained using official released code and hyperparameters were tuned to the best of our ability.
-TABLE I: Novel view synthesis performance on real world held-out validation data.
+All baselines were trained using official code and hyperparameters were tuned to the best of our ability.
+TABLE I: Novel view synthesis performance on real world held-out validation data (7 datasets).
 Note that ZSplat* is ZSplat with only the sonar loss activated for training.
-First, second, and third best are shown in red, orange, and yellow, respectively.
+First, second, and third best are shown in red, orange, and yellow.
 Sonar Image Synthesis
@@
 SonarSplat and ZSplat* both leverage Gaussian splatting, which offers significantly faster rendering speeds compared to neural rendering methods.
-ZSplat* is optimized for high rendering speed, which will be a focus of future work.
+ZSplat* is optimized for fast rendering, which will be a focus of future work.
+TABLE II: Ablation study results for the proposed acoustic reflectance parameter (VDR), azimuth streak modeling (ASM), and elevation sampling densification strategy (ESDS). Best in bold.
 SonarSplat Rendering Ablations
@@
 22.79 0.41 0.48
-TABLE II: Ablation study results for the proposed acoustic reflectance parameter (VDR), azimuth streak modeling (ASM), and elevation sampling densification strategy (ESDS). Best in bold.
 IV-D Ablation Studies
@@
 Note that if ASM is not modeled, rendered views will produce erroneous streaks (indicated with red arrows) instead of capturing the true streaks (indicated with blue arrows).
-Images are polar.
 Figure 7: Qualitative results for azimuth streak removal capabilities of SonarSplat. Locations of azimuth streaks in the captured images are indicated by blue arrows.
-TABLE III: Quantitative results for azimuth streak removal. We report the inverse coefficient of variation (ICV). Higher is better ($\uparrow$) and best is in bold.
+TABLE III: We report the inverse coefficient of variation (ICV) for azimuth streak removal. Best is in bold.
 De-Streaking Performance (ICV) $\uparrow$
@@
 The Azimuth Streak Modeling allows SonarSplat to qualitatively and quantitatively remove streaks effectively and improves the finer features in sonar images.
-Figure 8: Qualitative results from 3D reconstruction of SonarSplat compared to baselines. We inspect a concrete piling in a test tank. Ground truth is created using collected RGB images.
+Figure 8: Qualitative results from 3D reconstruction of SonarSplat compared to baselines.
+We inspect Concrete Piling in a test tank.
+Monohansett Shipwreck was surveyed in Lake Huron, Michigan.
+Ground truth is created using collected RGB images and LIDAR respectively.
+Readers are encouraged to zoom in.
 IV-F Real-World 3D Reconstruction
-3D reconstruction is crucial for inspection and visualization of underwater structures. In a test tank, we demonstrate inspection of a concrete piling, a structure commonly found in bridges and dams.
+3D reconstruction is crucial for inspection and visualization of underwater structures.
+In a test tank, we demonstrate inspection of Concrete Piling, a structure commonly found in bridges and dams.
+We also surveyed Monohansett Shipwreck, a wooden shipwreck at a depth of 18 feet in Lake Huron, Michigan, measuring 160 feet in length and 30 feet in width.
 We show that SonarSplat learns the geometric properties of the scene by converting the optimized splat into a mesh for visualization.
+We sample a set of points from the probability density functions described by the Gaussians, then query the opacities and use marching cubes to produce a mesh.
+Specifically, we used between 10-50 samples per Gaussian depending on the size of the scene and the density of the Gaussians.
 We compare to state-of-the-art sonar reconstruction methods: Neusis, DSC, and ZSplat* [13, 10, 19].
+Additionally, we ablate ESDS (denoted as “SonarSplat w/o ESDS”).
 Qualitative results are shown in Fig.
@@
 IV.
-Following [33, 6], we report Root Mean Squared (RMS) $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD).
+Following [10, 13], we report Root Mean Squared (RMS) $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD), which are both in meters.
 We produce meshes from all methods then randomly sample 30,000 points from the surface for 30 iterations to calculate RMS.
+We use ICP to align the predictions and ground truth meshes then crop the predictions to the bounding box of the ground truth.
 Note that neither DSC, Neusis, nor ZSplat* is able to reconstruct the object whereas SonarSplat is able to produce an accurate reconstruction of the piling.
-TABLE IV: Quantitative evaluation of 3D reconstruction performance on the piling using RMS $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD). Best in bold.
-Reconstruction Metrics
-Method CD-$l_{1}$downarrow HD $\downarrow$
-Neusis [33] 0.32 0.78
-DSC [6] 0.29 0.82
-ZSplat* [35] 2.93 4.88
-SonarSplat (ours) 0.14 0.57
+For Monohansett Shipwreck, we see that SonarSplat captures details like the wooden ribs of the shipwreck more accurately than baselines.
+TABLE IV: Quantitative evaluation of 3D reconstruction using RMS $l1$ Chamfer Distance (CD-$l1$) and Hausdorff Distance (HD). Units are meters. Best in bold.
+3D Reconstruction Performance
+Concrete Piling Monohansett Shipwreck
+Method CD-$l_{1}$ $\downarrow$ HD $\downarrow$ CD-$l_{1}$ $\downarrow$ HD $\downarrow$
+Neusis [10] 0.32 0.78 1.75 3.67
+DSC [13] 0.29 0.82 1.60 6.77
+ZSplat* [19] 2.93 4.88 3.06 8.30
+SonarSplat w/o ESDS 0.92 1.77 3.62 12.52
+SonarSplat (ours) 0.14 0.57 0.29 1.06
 V Conclusion & Future Work
@@
 We also introduce a novel densification strategy (ESDS) to enable SonarSplat to resolve elevation angle ambiguities and associated occlusions.
-From our extensive experiments on a diverse set of real-world data in test tank and river environments, we find that SonarSplat outperforms state-of-the-art methods in novel view synthesis (+3.2 dB PSNR) and 3D reconstruction (52% lower Chamfer Distance), and additionally allows for de-streaking of sonar images.
+From our extensive experiments on a diverse set of real-world data in test tank and lake environments, we find that SonarSplat outperforms state-of-the-art methods in novel view synthesis (+3.2 dB PSNR) and 3D reconstruction (77% lower Chamfer Distance), and additionally allows for de-streaking of sonar images.
 Future work will focus on leveraging SonarSplat for sonar data synthesis by randomizing scene parameters including acoustic reflectance and azimuth streaking probability.
 This can yield more diverse datasets for training sonar-based scene understanding algorithms, including multi-view target recognition algorithms [46, 47].
-We will also work to improve SonarSplat’s ability to perform dense 3D reconstruction by better modeling additional acoustic phenomena like speckle noise, Lambertian scattering, and multi-path reflections.
-Our choice of Gaussian splatting as a 3D scene representation also opens possibilities for real-time sensor fusion and SLAM [13, 23].
+We will also work to improve SonarSplat’s ability to perform dense 3D reconstruction by better modeling additional acoustic phenomena like speckle noise and multi-path reflections.
+Our choice of Gaussian splatting as a 3D scene representation also opens possibilities for real-time sensor fusion and SLAM [24].